### v1

# Non-local Hamilton–Jacobi-Bellman equations for the stochastic optimal control of path-dependent piecewise deterministic processes

Abstract.
We study the optimal control of path-dependent piecewise deterministic processes. An appropriate dynamic programming principle is established. We prove that the associated value function is the unique minimax solution of the corresponding non-local path-dependent Hamilton–Jacobi–Bellman equation. This is the first well-posedness result for nonsmooth solutions of fully nonlinear non-local path-dependent partial differential equations.
a University of Bologna, Department of Mathematics, Piazza di Porta San Donato 5, 40126 Bologna (Italy).
b University of Central Florida, Department of Mathematics, 4393 Andromeda Loop N Orlando, FL 32816 (USA).
1 E-mail: elena.bandini7@unibo.it.
2 E-mail: christian.keller@ucf.edu.
The research of the first named author was partially supported by the 2018 GNAMPA-INdAM project Controllo ottimo stocastico con osservazione parziale: metodo di randomizzazione ed equazioni di Hamilton-Jacobi-Bellman sullo spazio di Wasserstein and the 2024 GNAMPA-INdAM Problemi di controllo ottimo in dimensione infinita. The research of the second named author was supported in part by NSF-grant DMS-2106077.
Keywords: Path-dependent piecewise deterministic processes; non-local path-dependent HJB equations; stochastic optimal control.
AMS 2020: 45K05, 35D99, 49L20, 90C40, 93E20
1. Introduction
Piecewise deterministic Markov processes (PDMPs) have been popularized by Davis (see especially [Davis84] and [DavisBook]) and found many applications. PDMPs are pure jump processes; between consecutive jump times, they are deterministic and solve ordinary differential equations. We study a non-Markovian counterpart of PDMPs. More specifically, our processes, which we shall call path-dependent piecewise deterministic processes (PDPs), differ from (and extend) PDMPs inasmuch as all relevant data are allowed to depend on the (realized) process history as opposed to just the (realized) state as is the case for PDMPs. E.g., between consecutive jump times, path-dependent PDPs solve (delay) functional differential equations. Consequently, our theory broadens the realm of possible applications.111 General, not necessarily Markovian, PDPs are briefly mentioned in [Jacobsen]. We are not aware of any thorough treatment of path-dependent PDPs.
Our work focuses on stochastic optimal control problems that involve path-dependent PDPs and their related Hamilton–Jacobi–Bellman (HJB) equations. These equations are non-local and path-dependent. More precisely, our HJB equations are of the following form:
(1.1) $\begin{split}&\partial_{t}u(t,x)+\inf_{a\in A}\Big{\{}\lambda(t,x,a)\,\int_{{\mathbb{R}}^{d}}[u(t,{\mathbf{1}}_{[0,t)}\,x+{\mathbf{1}}_{[t,\infty)}e)-u(t,x)]\,Q(t,x,a,{\mathrm{d}}e)\\ &\qquad+\ell(t,x,a)+(f(t,x,a),\partial_{x}u(t,x))\Big{\}}=0,\quad(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d}).\end{split}$
Here, $D({\mathbb{R}}_{+},{\mathbb{R}}^{d})$ denotes the Skorokhod space. Also note that all terms in (1.1) are to be understood as non-anticipating, e.g., $u(t,x)=u(t,x_{\cdot\wedge t})$. Counterparts of (1.1) in the non-path-dependent case would be of the form
(1.2) $\begin{split}&\partial_{t}\tilde{u}(t,\tilde{x})+\inf_{a\in A}\{\tilde{\lambda}(t,\tilde{x},a)\,\int_{{\mathbb{R}}^{d}}[\tilde{u}(t,e)-\tilde{u}(t,\tilde{x})]\,\tilde{Q}(t,\tilde{x},a,{\mathrm{d}}e)\\ &\qquad+\tilde{\ell}(t,\tilde{x},a)+(\tilde{f}(t,\tilde{x},a),\partial_{\tilde{x}}\tilde{u}(t,\tilde{x}))\}=0,\quad(t,\tilde{x})\in[0,T)\times{\mathbb{R}}^{d}.\end{split}$
As a rule, classical solutions to such HJB equations rarely exist and theories of non-smooth solutions such as viscosity or minimax solutions are needed. In the path-dependent case, the only relevant work for non-local equations is [Keller16SPA], where viscosity solutions for semilinear non-local path-dependent partial differential equations are studied. But there have been no works on nonsmooth solutions for fully nonlinear non-local path-dependent partial differential equations such as (1.1). This is in stark contrast to the Markovian case: Partial differential equations associated to PDMPs, i.e., variations of (1.2), were studied in [Vermes85Stoch], [Soner86SICON-II], [Ye94] [DempsterYe95AAP], [DavisFarid], [ForwickEtAl04], [Goreac12], and [BandiniThieullen20AMO] to mention but a few works.
1.1. A simplified presentation of the control problem
Given initial conditions $(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})$, our stochastic optimal control problem is to minimize a cost functional
$J(t,x,\alpha):={{\mathbb{E}}}\left[\int_{t}^{T}\ell(s,X_{\cdot\wedge s}^{t,x,\alpha},\alpha(s))\,{\mathrm{d}}s+h(X_{\cdot\wedge T}^{t,x,\alpha})\right]$
over an appropriate class of controls $\alpha$. The controlled stochastic process $X^{t,x,\alpha}$ is a path-dependent piecewise deterministic process with jump times $T_{n}^{t,x,\alpha}$, $n\in{\mathbb{N}}$, and post-jump locations $E_{n}^{t,x,\alpha}$, $n\in{\mathbb{N}}$, such that the the following holds:
$X^{t,x,\alpha}(s)=x(s)\quad\text{for each $s\in[0,t]$},$
$T_{0}^{t,x,\alpha}=t,\quad\text{$T_{1}^{t,x,\alpha}$ is the first jump time of $X^{t,x,\alpha}$, etc.}$.
$E_{0}^{t,x,\alpha}=x(t),\quad\text{$E_{1}^{t,x,\alpha}$ is the first post-jump location of $X^{t,x,\alpha}$, }$
i.e., $E_{1}^{t,x,\alpha}=X^{t,x,\alpha}(T_{1}^{t,x,\alpha})$, etc., and
$\frac{d}{ds}X^{t,x,\alpha}(s)=f(s,X_{\cdot\wedge s}^{t,x,\alpha},\alpha(s))\text{ for a.e.~{}$s\in(T_{n}^{t,x,\alpha},T_{n+1}^{t,x,\alpha})$.}$
The probability distributions of the jump times and post-jump locations are specified via a survival function $F^{t,x,\alpha}$ and a random measure $Q(\cdot,{\mathrm{d}}e)$ on ${\mathbb{R}}^{d}$, e.g.,
${{\mathbb{P}}}(T^{t,x,\alpha}_{1}>s)$ $=F^{t,x,\alpha}(s)\text{ for every $s>T$ and}$
${{\mathbb{P}}}(E^{t,x,\alpha}_{1}\in B|T^{t,x,\alpha}_{1})$ $=Q(T^{t,x,\alpha}_{1},\phi_{\cdot\wedge T^{t,x,\alpha}_{1}}^{t,x,\alpha},\alpha(T^{t,x,\alpha}_{1}),B)\text{ for each Borel set $B\subset{\mathbb{R}}^{d}$,}$
where $\phi^{t,x,\alpha}$ is [^2] Subject to our data satisfy appropriate conditions. (!) the unique solution of $\frac{d}{ds}\phi^{t,x,\alpha}(s)=f(s,\phi_{\cdot\wedge s}^{t,x,\alpha},\alpha(s))$ on $(t,\infty)$ with initial condition $\phi^{t,x,\alpha}(s)=x(s)$ on $[0,t]$, i.e., $\phi^{t,x,\alpha}$ and $X^{t,x,\alpha}$ coincide before the first jump time.
The class of controls: Differences between the Markovian and the path-dependent cases. In both cases, controls are open-loop between jump times. In the Markovian case, admissible controls are of the form [^3] For readability’s sake, we omit the superscript $\phantom{}{}^{t,x,\alpha}$ or part of it.
$\tilde{\alpha}(t)=\tilde{\alpha}_{n}(t|T_{n},E_{n}),\quad T_{n}\leq t<T_{n+1}.$
This structure leads naturally to an equivalent (discrete-time) Markov decision model (see, e.g., [DavisBook]). [^4] Such an equivalency is of major importance in the theory of PDMPs and suitable counterparts are crucial in this work as well. For details, see section 4.5 and appendix A. In the path-dependent case, controls of the form
(1.3) $\alpha(t)=\alpha_{n}(t|T_{n},\{X^{\alpha}(s)\}_{0\leq s\leq{T_{n}}}),\quad T_{n}\leq t<T_{n+1},$
where $X^{\alpha}$ is our controlled piecewise deterministic process, ought to be at least a subset of all admissible controls. To establish a correspondence to an equivalent discrete-time model, we would need to express the history $\{X^{\alpha}(s)\}_{0\leq s\leq t}$ as an appropriate function of finitely many relevant random variables such as the past jump times $T_{1}$, $\ldots$, $T_{n}$, and past jump locations $E_{1}$, $\ldots$, $E_{n}$. However, to determine $X^{\alpha}$ on $[0,t]$, we additionally [^5] Unlike in the Markovian case, where only $X^{\alpha}(T_{n})=E_{n}$ is needed. need to know the control history, i.e., all (realized) open-loop controls $\alpha_{0}$, $\ldots$, $\alpha_{n-1}$. [^6] This will lead to larger class of admissible controls compared to the ones described in (1.3). What makes matters slightly complicated is that each such (realized) open-loop control $\alpha_{k}$ also depends on $T_{1}$, $\ldots$, $T_{k}$, $E_{1}$, $\ldots$, $E_{k}$, and $\alpha_{0}$, $\ldots$, $\alpha_{k-1}$. To make the situation more tractable, we will use randomized policies (see Definition 4.7). Finally note that such a seemingly complicated setting for our controls is needed in the proof of the dynamic programming principle (Theorem 4.10), which is crucial in establishing that solutions of the HJB equation (1.1) coincide with the value function of our stochastic optimal control problem.
1.2. Our approach for the non-local path-dependent HJB equation (1.1)
We follow the methodology by Davis and Farid in [DavisFarid], where existence and uniqueness of viscosity solutions for HJB equations related to PDMPs are established via a fixed-point argument. However, our situation in the path-dependent case leads to several additional obstacles. We continue by describing some of those difficulties and how they will be circumvented. First note that main ingredients of the Davis–Farid methodology are existence and uniqueness results of certain local HJB equations, in our case, these HJB equations are path-dependent and of the form
(1.4) $\begin{split}\partial_{t}u(t,x)+F_{\psi}(t,x,u(t,x),\partial_{x}u(t,x))=0,\quad(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d}),\end{split}$
where $\psi:[0,T]\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})\to{\mathbb{R}}$ and $F_{\psi}=F_{\psi}(t,x,y,z)$ is of the form
(1.5) $\begin{split}&F_{\psi}(t,x,y,z)=\inf_{a\in A}\{\lambda(t,x,a)\int_{{\mathbb{R}}^{d}}[\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)-y]\,Q(t,x,a,{\mathrm{d}}e)\\ &\quad+\ell(t,x,a)+(f(t,x,a),z)\},\quad(t,x,y,z)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})\times{\mathbb{R}}\times{\mathbb{R}}^{d}.\end{split}$
Roughly speaking, a function $u:[0,T]\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})\to{\mathbb{R}}$ is said to be a solution of (1.1) in some sense if $u$ is a solution of (1.4) with $\psi\equiv u$ in the same sense.
One of our obstacles in the path-dependent case is a possible lack of regularity of the term $\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)$ in (1.5). Even if $\psi$ is continuous, the map $t\mapsto\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)$ might be not continuous (see appendix C) and therefore we can expect the Hamiltonian $F_{\psi}$ to be at most measurable with respect to $t$. In our companion paper [BK1], we develop a minimax solution theory for path-dependent Hamilton–Jacobi equations with time-measurable coefficients and also with “$u$-dependence” of the Hamiltonian. There seems to have been no treatment of any of those two aspects in the literature (see [GLP21AMO] for an overview) and thus the results in [BK1] ought to be of independent interest besides the current matter.
A second obstacle is the following. A standard proof of a comparison principle between appropriate sub- and supersolutions of (1.4) (as done in [BK1]) requires at least some regularity of $F_{\psi}$ with respect to its $x$-component. Typically, some type of uniform continuity is required. For simplicity (even though it might be more difficult to prove), let us require Lipschitz continuity, i.e., we would need [^7] We show only the term that causes most trouble.
$\left|{\int_{{\mathbb{R}}^{d}}[\psi(t,{\mathbf{1}}_{[0,t)}x_{1}+{\mathbf{1}}_{[t,\infty)}e)-\psi(t,{\mathbf{1}}_{[0,t)}x_{2}+{\mathbf{1}}_{[t,\infty)}e)\,Q(t,x_{1},a,{\mathrm{d}}e)}\right|$
$\qquad\qquad\leq L_{\psi}\sup_{0\leq s\leq t}\left|{x_{1}(s)-x_{2}(s)}\right|,$
which, in the Markovian or state-dependent case, is trivially satisfied, as
$\psi(t,{\mathbf{1}}_{[0,t)}x_{1}+{\mathbf{1}}_{[t,\infty)}e)-\psi(t,{\mathbf{1}}_{[0,t)}x_{2}+{\mathbf{1}}_{[t,\infty)}e)=\tilde{\psi}(t,e)-\tilde{\psi}(t,e)=0$
for some function $\tilde{\psi}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ with $\tilde{\psi}(t,x(t))=\psi(t,x)$. The slightly involved proof of Lemma 5.7 deals with this issue by establishing Lipschitz continuity for our value function $V$, which yields the desired regularity at least for $F_{\psi}$ with $\psi=V$.
1.3. Organization of the rest of the paper
Section 2 introduces frequently used notation. In section 3, two canonical sample spaces are specified, the path space $\Omega$ (the Skorokhod space, previously denoted by $D({\mathbb{R}}_{+},{\mathbb{R}}^{d})$) and the canonical space of our controlled marked point processes. Section 4 provides a precise description of our stochastic optimal control problem including assumptions for the data as well as statement and proof of the dynamic programming principle. Section 5 contains the complete treatment of our non-local path-dependent HJB equation: The fixed-point approach in the spirit of Davis and Farid, appropriate notions of minimax solutions, and, ultimately, existence, uniqueness, and a comparison principle for (1.1) together with relevant terminal conditions. The appendices include technical as well as fundamental material. Appendix A is relevant in its own right (independent from its applications in this paper): It connects a fairly general path-dependent discrete-time decision model to a standard discrete-time decision model. Thereby, the Bellman equation for the path-dependent model is derived. Appendix B contains technical proofs. Finally, appendix C provides results on a possible lack of regularity of certain functionals on Skorokhod space.
2. Notation
Let ${\mathbb{N}}$ be the set of all strictly positive integers and ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$. Fix $d\in{\mathbb{N}}$ and $T\in(0,\infty)$. Let ${\mathbb{R}}_{+}=[0,\infty)$. Let $E={\mathbb{R}}^{d}$ with cemetery state $\Delta\not\in E$ Given $(s,x)\in{\mathbb{R}}_{+}\times D({\mathbb{R}}_{+},E)$ and $e\in E$, define
(2.1) $x\otimes_{s}e\quad\text{by\,\,\,\, $(x\otimes_{s}e)(t)=\begin{cases}x(t),\,\textup{if}\,\,t\in[0,s)\\ e,\,\textup{if}\,\,t\in[s,\infty).\end{cases}$}$
For measurability properties in $(x,t,e)$, see Theorem 96, 146-IV, in [DMprobPot].
Remark 2.1.
Clearly, $x\otimes_{s}e=x(\cdot\wedge s)\otimes_{s}e$. This identity will simplify notation.
Given a topological space $S$, we denote by $\mathcal{B}(S)$ its Borel $\sigma$-field, by $\mathscr{P}(S)$ the set of all probability measures on $\mathcal{B}(S)$, and by $B(S)$ the set of all bounded Borel-measurable functions from $S$ to ${\mathbb{R}}$. Moreover, we write $C(S)$, $\mathrm{USC}(S)$, and $\mathrm{LSC}(S)$ for the sets of all continuous, upper semicontinuous, and lower semicontinuous functions from $S$ to ${\mathbb{R}}$, resp. Given another topological space $\tilde{S}$, we write $C(S,\tilde{S})$ for the set of all continuous functions from $S$ to $\tilde{S}$.
We frequently use the notation ${\mathbf{1}}_{S}$ for indicator functions of some set $S$, i.e., ${\mathbf{1}}_{S}(t)=1$ if $t\in S$ and ${\mathbf{1}}_{S}(t)=0$ if $t\not\in S$, and the notation $\delta_{r}$ for a Dirac measure concentrated at some point $r$; the domains of ${\mathbf{1}}_{S}$ and of $\delta_{r}$ will be clear from context. We also always use the convention $\infty\cdot 0=0$.
3. The canonical setup
3.1. The canonical path space
Let $\Omega=D({\mathbb{R}}_{+},E)$. The canonical process $X=(X_{t})_{t\geq 0}$ on $\Omega$ is defined by $X_{t}(\omega)=\omega(t)$ for each $(t,{\omega})\in{\mathbb{R}}_{+}\times\Omega$. The random measure associated to the jumps of $X$ is denoted by $\mu^{X}$, i.e.,
$\mu^{X}({\mathrm{d}}t\,{\mathrm{d}}e)=\sum_{r}{\mathbf{1}}_{\{\Delta X_{r}\neq 0\}}\,\delta_{(r,X_{r})}({\mathrm{d}}t\,{\mathrm{d}}e).$
Here, $\Delta X_{r}=X_{r}-X_{r-}$. Let ${\mathbb{F}}^{0}=\{{\mathcal{F}}^{0}_{t}\}_{t\geq 0}$ be the filtration generated by $X$ and equip $\Omega$ with the $\sigma$-field ${\mathcal{F}}^{0}:=\vee_{t\geq 0}{\mathcal{F}}^{0}_{t}$. We also use the right-continuous filtration ${\mathbb{F}}=\{{\mathcal{F}}_{t}\}_{t\geq 0}$ defined by ${\mathcal{F}}_{t}:=\cap_{{\varepsilon}>0}{\mathcal{F}}^{0}_{t+{\varepsilon}}$ and the $\sigma$-field ${\mathcal{F}}:=\vee_{t\geq 0}{\mathcal{F}}_{t}$.
We consider non-empty subsets of +×${\mathbb{R}}_{+}\times\Omega$ to be equipped with the pseudo-metric
(3.1) $((t,x),(s,\tilde{x}))\mapsto\left|{t-s}\right|+\sup_{0\leq r<\infty}\left|{x(r\wedge t)-\tilde{x}(r\wedge s)}\right|.$
3.2. The canonical space of controlled marked point processes
First, consider a non-empty Borel subset $A$ of a Polish space and fix an element $a^{\circ}\in A$. The set $A$ is our control action space. Next, define a set of open-loop controls
[…]
with cemetery state $\Delta^{\prime}\not\in\mathcal{A}$.
Remark 3.1.
Note that $\mathcal{A}$ is a Borel space when it is equipped with $\mathcal{B}(\mathcal{A})$, the smallest $\sigma$-algebra under which all functions $\alpha\mapsto\int_{0}^{\infty}{\mathrm{e}}^{-t}\eta(t,\alpha(t))\,{\mathrm{d}}t$, $\mathcal{A}\to{\mathbb{R}}$, $\eta\in B({\mathbb{R}}_{+}\times A)$, are measurable (Lemma 1 in [Yushkevich80]).
Our canonical space $\check{\Omega}$ is defined as the set of all sequences $\check{\omega}=(t_{n},e_{n},\alpha_{n})_{n\in{\mathbb{N}}_{0}}$ for which the following holds (cf. Remarque III.3.43 in [JacodBook]):
(i) $(t_{n},e_{n},\alpha_{n})\in\left((0,\infty)\times E\times\mathcal{A}\right)\cup\{(\infty,\Delta,\Delta^{\prime})\}$.
(ii) $t_{n}\leq t_{n+1}$.
(iii) $t_{n}<t_{n+1}$ unless $t_{n}=\infty$.
For each $n\in{\mathbb{N}}_{0}$, define canonical mappings ${T}_{n}:\check{\Omega}\to(0,\infty]$, ${E}_{n}:\check{\Omega}\to E\cup\{\Delta\}$, and $\mathcal{A}_{n}:\check{\Omega}\to\mathcal{A}\cup\{\Delta^{\prime}\}$ by
${T}_{n}(\check{\omega})=t_{n},\quad{E}_{n}(\check{\omega})=e_{n},\quad\mathcal{A}_{n}(\check{\omega})=\alpha_{n},\qquad\check{\omega}=(t_{n},e_{n},\alpha_{n})_{n\in{\mathbb{N}}_{0}}\in\check{\Omega}.$
Consider the corresponding random measure
$\check{\mu}({\mathrm{d}}t\,{\mathrm{d}}e\,{\mathrm{d}}\alpha)=\sum_{n\geq 0}{\mathbf{1}}_{\{T_{n}<\infty\}}\,\delta_{({T}_{n},{E}_{n},\mathcal{A}_{n})}({\mathrm{d}}t\,{\mathrm{d}}e\,{\mathrm{d}}\alpha),$
let $\check{{\mathbb{F}}}=\{\check{{\mathcal{F}}}_{t}\}_{t\geq 0}$ be the filtration [^8] This is a raw filtration. It is emphasized in [Jacobsen] that completed filtrations should not be used. generated by $\check{\mu}$, and equip $\check{\Omega}$ with the $\sigma$-field ˇ$\check{{\mathcal{F}}}:=\vee_{t\geq 0}\check{{\mathcal{F}}}_{t}$.
Remark 3.2.
The filtration $\check{{\mathbb{F}}}$ is right-continuous (Theorem T25, p. 304, in [BremaudBook]).
For each $n\in{\mathbb{N}}$, we shall also need the space [^9] Our numbering is consistent with the one of the marginals $r_{N}$ in chapters 8 and 9 of [BS].
$\check{\Omega}_{n}$ :=$:=\{\check{\omega}_{n}\in(({\mathbb{R}}_{+}\times E\times\mathcal{A})\cup\{(\infty,\Delta,\Delta^{\prime})\})^{n}:$
$\qquad\qquad\exists(t_{j},e_{j},\alpha_{j})_{j\in{\mathbb{N}}_{0}}\in\check{\Omega}\text{ with }\check{\omega}_{n}=(t_{j},e_{j},\alpha_{j})_{j=0}^{n-1}\}$
with corresponding filtration $\check{{\mathbb{F}}}_{n}=\{\check{{\mathcal{F}}}_{n,t}\}_{t\geq 0}\}$ and $\sigma$-field ˇ$\check{{\mathcal{F}}}_{n,\infty}:=\vee_{t\geq 0}\check{{\mathcal{F}}}_{n,t}$. With slight abuse of notation, we write $T_{j}$, $E_{j}$, and $\mathcal{A}_{j}$ for the canonical mappings on $\check{\Omega}_{n}$ and do not indicate their domains (i.e., $\check{\Omega}_{n}$ vs. $\check{\Omega}$) if there is no danger of confusion.
4. Optimal control of path-dependent PDPs
4.1. The data
First, consider mappings
$f:{\mathbb{R}}_{+}\times\Omega\times A\to E,\quad\lambda:{\mathbb{R}}_{+}\times\Omega\times A\to{\mathbb{R}}_{+},\quad Q:{\mathbb{R}}_{+}\times\Omega\times A\times\mathcal{B}(E)\to{\mathbb{R}}_{+},$
which are the characteristics for our PDP. Next, consider mappings
:+×$\ell:{\mathbb{R}}_{+}\times\Omega\times A\to{\mathbb{R}}_{+}\qquad\text{and}\qquad h:\Omega\to{\mathbb{R}}_{+},$
which will serve as running and terminal costs.
The following standing assumptions are in force throughout the rest of this paper.
Assumption 4.1.
Let $f$, $\lambda$, $\ell$, and $h$ satisfy the following conditions (the same as in section 7 in [BK1]).
(i) For a.e. $t\in{\mathbb{R}}_{+}$, the map $(x,a)\mapsto(f,\lambda,\ell)(t,x,a)$, $\Omega\times A\to E\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}$, is continuous, where $\Omega$ is considered here to be equipped with the seminorm $x\mapsto\sup_{s\leq t}\left|{x(s\wedge t)}\right|$.
(ii) For every $(x,a)\in\Omega\times A$, the map $t\mapsto(f,\lambda,\ell)(t,x,a)$, ${\mathbb{R}}_{+}\to E\times{\mathbb{R}}_{+}\times{\mathbb{R}}_{+}$, is Borel-measurable.
(iii) There are constants $C_{f}$, $C_{\lambda}\geq 0$, $L_{f}\geq 0$, such that, for every $(s,x)\in{\mathbb{R}}_{+}\times\Omega$,
$\sup_{a\in A}\left|{f(s,x,a))}\right|\leq C_{f}(1+\sup_{t\leq s}\left|{x(t)}\right|),\quad\sup_{a\in A}\lambda(t,x,a)\leq C_{\lambda},\qquad\text{and}$
$\sup_{a\in A}\ell(s,x,a)+h(x)\leq C_{f}.$
(iv) There is a constant $L_{f}\geq 0$ such that, for every $(s,x,\tilde{x})\in{\mathbb{R}}_{+}\times\Omega\times\Omega$,
$\sup_{a\in A}\left[\left|{f(s,x,a)-f(s,\tilde{x},a)}\right|+\left|{\ell(s,x,a)-\ell(s,\tilde{x},a)}\right|+\left|{\lambda(s,x,a)-\lambda(s,\tilde{x},a)}\right|\right]$
$\qquad\leq L_{f}\sup_{t\in[0,s]}\left|{x(t)-\tilde{x}(t)}\right|\qquad\text{and}$
$\left|{h(x)-h(\tilde{x})}\right|\leq L_{f}\sup_{t\leq T}\left|{x(t)-\tilde{x}(t)}\right|.$
Assumption 4.2.
Let $Q$ satisfy the following conditions.
(i) For every $B\in\mathcal{B}(E)$, $(s,x,a)\mapsto Q(s,x,a,B)$ is Borel-measurable and non-anticipating, i.e., $Q(s,x,a,B)=Q(s,x(\cdot\wedge s),a,B)$.101010 By Theorem 97, 147-IV, in [DMprobPot], the maps $(s,x)\mapsto Q(s,x,a,B)$, $(a,B)\in A\times\mathcal{B}(E)$, are ${\mathbb{F}}^{0}$-optional.
(ii) For every $s\in{\mathbb{R}}_{+}$, $x\in\Omega$, and $a\in A$, $Q(s,x,a,\cdot)\in\mathscr{P}(E)$.
(iii) For every $s\in{\mathbb{R}}_{+}$, $x\in\Omega$, and $a\in A$, $Q(s,x,a,\{x(s)\})=0$.
(iv) There is a constant $L_{Q}\geq 0$, such that, for all $\psi\in C_{b}([0,T]\times\Omega)$ and $L_{\psi}\geq 0$, the statement
$\left|{\psi(s,x)-\psi(s,\tilde{x})}\right|\leq L_{\psi}{\color[rgb]{0,0,0}\sup_{t\in[0,s]}\left|{x(t)-\tilde{x}(t)}\right|}\text{ for all $(s,x,\tilde{x})\in[0,T]\times\Omega\times\Omega$}$
implies
$\sup_{a\in A}\Big{|}\int_{{\mathbb{R}}^{d}}\psi(s,x\otimes_{s}e)\,Q(s,x,a,{\mathrm{d}}e)-\int_{{\mathbb{R}}^{d}}\psi(s,x\otimes_{s}e)\,Q(s,\tilde{x},a,{\mathrm{d}}e)\Big{|}$
$\qquad\leq L_{\psi}\,L_{Q}\,\sup_{t\in[0,s]}\left|{x(t)-\tilde{x}(t)}\right|\text{ for all $(s,x,\tilde{x},e)\in[0,T]\times\Omega\times\Omega\times E$.}$
Remark 4.3.
By Proposition 7.26 in [BS], $Q$ is a Borel-measurable stochastic kernel on $E$ given +×${\mathbb{R}}_{+}\times\Omega\times A$.
4.2. The flow and related notation
Given $(s,x,\alpha)\in{\mathbb{R}}_{+}\times\Omega\in\mathcal{A}$, denote by $\phi^{s,x,\alpha}=\phi$ the solution of
(4.1) $\begin{split}\phi^{\prime}(t)&=f(t,\phi,\alpha(t))\text{ a.e.~{}on $(s,\infty)$,}\\ \phi(t)&=x(t)\text{ on $[0,s]$.}\end{split}$
Establishing existence and uniqueness for the initial value problem (4.1) is standard given Assumption 4.1. Note that $\phi^{s,x,\alpha}\in\Omega$ and $\phi^{s,x,\alpha}|_{[s,\infty)}\in C([s,\infty),E)$.
The following result is implicitly needed throughout most of what follows. Its proof can be found in the appendix, section B.1.
Lemma 4.4.
The map $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$.
We shall also make use of the survival function $F^{s,x,\alpha}$ defined by
$F^{s,x,\alpha}(t)$ $:={\mathbf{1}}_{[s,\infty)}(t)\,\exp\Big{(}-\int_{s}^{t}\lambda(r,\phi^{s,x,\alpha},\alpha(r))\,{\mathrm{d}}r\Big{)}+{\mathbf{1}}_{[0,s)}(t),\quad t\in{\mathbb{R}}_{+}.$
For each $(s,x,\alpha)\in{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}$ and $t\in[s,\infty)$, also define
$\lambda^{s,x,\alpha}(t)$ $:=\lambda(t,\phi^{s,x,\alpha},\alpha(t)),$
$\chi^{s,x,\alpha}(t)$ $:=\exp\left(-\int_{s}^{t}\lambda^{s,x,\alpha}(r)\,{\mathrm{d}}r\right),$
$\ell^{s,x,\alpha}(t)$ $:=\ell(t,\phi^{s,x,\alpha},\alpha(t)),$
$Q^{s,x,\alpha}(t,{\mathrm{d}}e)$ $:=Q(t,\phi^{s,x,\alpha},\alpha(t),{\mathrm{d}}e).$
If more convenient, then we write
$Q(t,{\mathrm{d}}e|s,x,\alpha):=Q^{s,x,\alpha}(t,{\mathrm{d}}e),\text{ etc.}$.
Remark 4.5.
Following Lemma 4.4 and the arguments in the proofs in section B.1,111111 E.g., assume first that $\lambda=\lambda(t,x,a)$ is of the form $\lambda_{1}(t)\,\lambda_{2}(x(\cdot\wedge t))\,\lambda_{3}(a)$, where $\lambda_{1}$ is (Borel) measurable and $\lambda_{2}$ as well as $\lambda_{3}$ are continuous. Then, by Lemma 4.4, $(s,x,\alpha,t)\mapsto\lambda(t,\phi^{s,x,\alpha},\alpha(t))$ is clearly $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})\otimes\mathcal{B}({\mathbb{R}}_{+})$-measurable. It remains to apply a monotone-class argument. one can deduce that the mappings
$(s,x,\alpha,t)\mapsto\lambda^{s,x,\alpha}(t),{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\times{\mathbb{R}}_{+}\to{\mathbb{R}}_{+},$
$(s,x,\alpha,t)\mapsto\chi^{s,x,\alpha}(t),{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\times{\mathbb{R}}_{+}\to{\mathbb{R}}_{+},$
$(s,x,\alpha,t)\mapsto\ell^{s,x,\alpha}(t),{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\times{\mathbb{R}}_{+}\to{\mathbb{R}}_{+},$
$(s,x,\alpha,t)\mapsto Q^{s,x,\alpha}(t,B),{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\times{\mathbb{R}}_{+}\to{\mathbb{R}}_{+},B\in\mathcal{B}(E),$
are $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})\otimes\mathcal{B}({\mathbb{R}}_{+})$-measurable.
4.3. Path-dependent controlled PDPs on $\check{\Omega}$
Let $(s,x)\in(0,\infty)\times\Omega$. Define the process $X^{s,x}:{\mathbb{R}}_{+}\times\check{\Omega}\to E\cup\{\Delta\}$ recursively by
(4.2) $X^{s,x}(t,\check{\omega}):=\begin{cases}\Delta,&t_{0}=\infty,\\ \phi^{s,x,\alpha_{0}}(t),&0\leq t<t_{1}\text{ and }t_{0}<\infty,\\ \phi^{t_{n},X^{s,x}(\check{\omega})\otimes_{t_{n}}\,e_{n},\alpha_{n}}(t),&t_{n}\leq t<t_{n+1}\text{ and }t_{0}<\infty,\end{cases}$
for each $\check{\omega}=(t_{n},e_{n},\alpha_{n})_{n\in{\mathbb{N}}_{0}}\in\check{\Omega}$. Keep in mind that, by Remark 2.1, the $X^{s,x}$-term on the right-hand side of (4.2) depends only on its values on $[0,t_{n}]$, i.e., our definition of $X^{s,x}$ is not circular. Note that ${X}^{s,x}(\check{\omega})\in\Omega$ for each $\check{\omega}\in\check{\Omega}$ with $t_{0}<\infty$. Also note that in (4.2), $t_{n+1}=\infty$ for some $n\in{\mathbb{N}}_{0}$ is possible and if, in addition, $t_{n}<\infty$, then $X^{s,x}(t,\check{\omega})=\phi^{t_{n},X^{s,x}(\check{\omega})\otimes_{t_{n}}\,e_{n},\alpha_{n}}(t)$ for all $t\in[t_{n},\infty)$. In particular, $X^{s,x}(t,(t_{0},e_{0},\alpha_{0};\infty,\Delta,\Delta^{\prime};\infty,\Delta,\Delta^{\prime};\ldots))=\phi^{s,x}(t)$ if $t_{0}<\infty$.
Remark 4.6.
The process $X^{s,x}$ is $\check{{\mathbb{F}}}$-adapted. This can be shown via mathematical induction together with Proposition 4.2.1 (b) (iii) in [Jacobsen], Theorem 96 (d), 146-IV, in [DMprobPot], and Lemma 4.4. Since in addition $X^{s,x}$ has càdlàg paths, it is also $\check{{\mathbb{F}}}$-progressively measurable.
We also consider the following “restrictions” of $X^{s,x}$. Given $n\in{\mathbb{N}}_{0}$, define $X^{s,x}_{n}:{\mathbb{R}}_{+}\times\check{\Omega}_{n+1}\to E\cup\{\Delta\}$ by
(4.3) $X^{s,x}_{n}(t,\check{\omega}_{n+1}):=X^{s,x}(t,(\check{\omega}_{n+1};\infty,\Delta,\Delta^{\prime};\infty,\Delta,\Delta^{\prime};\ldots)),$
where $\check{\omega}_{n+1}=(t_{j},e_{j},\alpha_{j})_{j=0}^{n}$.
Definition 4.7.
Denote by $\mathbb{A}^{\prime}$ the space of all policies $\mathbf{a}=(\mathbf{a}_{0},\mathbf{a}_{1},\ldots)$ such that each $\mathbf{a}_{n}=\mathbf{a}_{n}({\mathrm{d}}\alpha_{n}|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n-1})$, $n\in{\mathbb{N}}_{0}$, is a universally measurable121212 Recall from Definition 7.18 in [BS] that a subset $\underline{S}$ of a Borel space $S$ is universally measurable if, for every ${{\mathbb{P}}}\in\mathscr{P}(S)$, the set $\underline{S}$ belongs to the completion of $\mathcal{B}(S)$ with respect to ${{\mathbb{P}}}$. stochastic kernel on $\mathcal{A}\cup\{\Delta^{\prime}\}$ given ((+×$(({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\})^{n+1}\times(\mathcal{A}\cup\{\Delta^{\prime}\})^{n}$.
Denote by $\mathbb{A}$ the space of all non-randomized policies $\mathbf{a}=(\mathbf{a}_{0},\mathbf{a}_{1},\ldots)\in\mathbb{A}^{\prime}$, i.e., for each $n\in{\mathbb{N}}_{0}$, $(t_{0},e_{0})$, $\ldots$, $(t_{n},e_{n})\in({\mathbb{R}}_{+}\times E)\cup\{\infty,\Delta\}$, and $\alpha_{0}$, $\ldots$, $\alpha_{n-1}\in\mathcal{A}\cup\{\Delta^{\prime}\}$, $\mathbf{a}_{n}({\mathrm{d}}\alpha_{n}|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n-1})$ is a Dirac measure.
Remark 4.8.
Our (randomized) policies should not be confused with relaxed controls, which, in our context, would be functions from $[0,T]$ to $\mathscr{P}(A)$.
We want to stress that our use of randomized policies is mainly done for the sake of simplicity. While the setting appears to become more complex compared to the use of path-dependent feedback controls, the notation actually becomes easier and the embedding of our continuous-time model into a discrete-time model becomes straight-forward (see also Remark A.2).
Definition 4.9.
Let $(s,x)\in(0,T)\times\Omega$ and $\mathbf{a}\in\mathbb{A}^{\prime}$. Construct a probability measure ˇ$\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}$ on $(\check{\Omega},\check{{\mathcal{F}}})$ with corresponding expected value $\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}$ recursively by defining its marginals ˇ$\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}_{n}$ on $(\check{\Omega}_{n},\check{{\mathcal{F}}}_{n,\infty})$ with corresponding expected values $\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{n}$, $n\in{\mathbb{N}}$, as follows (cf. section 3.2 of [Jacobsen]):
(4.4) $\begin{split}&\check{{{\mathbb{E}}}}_{1}^{s,x,\mathbf{a}}[\psi_{0}(T_{0},E_{0},\mathcal{A}_{0})]:=\int_{0}^{\infty}\int_{E}\int_{\mathcal{A}}\psi_{0}(t_{0},e_{0},\alpha_{0})\,\mathbf{a}_{0}({\mathrm{d}}\alpha_{0})\,\delta_{x(s)}({\mathrm{d}}e_{0})\,\delta_{s}({\mathrm{d}}t_{0}),\\ &\check{{{\mathbb{E}}}}_{2}^{s,x,\mathbf{a}}[{\mathbf{1}}_{\{T_{1}<\infty\}}\psi_{1}(T_{0},E_{0},\mathcal{A}_{0};T_{1},E_{1},\mathcal{A}_{1})]\\ &\,:=-\int_{0}^{\infty}\int_{E}\int_{\mathcal{A}}\int_{0}^{\infty}\int_{E}\int_{\mathcal{A}}\psi_{1}(t_{0},e_{0},\alpha_{0};t_{1},e_{1},\alpha_{1})\\ &\,\qquad\qquad\qquad\mathbf{a}_{1}({\mathrm{d}}\alpha_{1}|(t_{0},e_{0}),(t_{1},e_{1});\alpha_{0})\\ &\,\qquad\qquad\qquad Q^{s,x,\alpha_{0}}(t_{1},{\mathrm{d}}e_{1})\,{\mathrm{d}}F^{s,x,\alpha_{0}}(t_{1})\,\mathbf{a}_{0}({\mathrm{d}}\alpha_{0})\,\delta_{x(s)}({\mathrm{d}}e_{0})\,\delta_{s}({\mathrm{d}}t_{0}),\\ &\check{{{\mathbb{E}}}}_{n+2}^{s,x,\mathbf{a}}[{\mathbf{1}}_{\{T_{n+1}<\infty\}}\psi_{n+1}(T_{0},E_{0},\mathcal{A}_{0};\ldots;T_{n+1},E_{n+1},\mathcal{A}_{n+1})]\\ &\,:=-\int_{\check{\Omega}_{n+1}}\Biggl{\{}\int_{0}^{\infty}\int_{E}\int_{\mathcal{A}}\psi_{n+1}(\check{\omega}_{n+1};t_{n+1},e_{n+1},\alpha_{n+1})\\ &\qquad\qquad\mathbf{a}_{n+1}({\mathrm{d}}\alpha_{n+1}|(t_{0},e_{0}),\ldots,(t_{n+1},e_{n+1});\alpha_{0},\ldots,\alpha_{n})\\ &\qquad\qquad Q(t_{n+1},{\mathrm{d}}e_{n+1}|t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n})\,{\mathrm{d}}F(t_{n+1}|t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n})\Biggr{\}}\\ &\qquad\qquad\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}_{n+1}({\mathrm{d}}\check{\omega}_{n+1}),\end{split}$
where $\psi_{n}:({\mathbb{R}}_{+}\times E)^{n+1}\times\mathcal{A}^{n+1}\to{\mathbb{R}}$, $n\in{\mathbb{N}}_{0}$, are Borel-measurable and bounded and the notation $\check{\omega}_{n+1}=(t_{j},e_{j},\alpha_{j})_{j=0}^{n}$ is used.
4.4. The optimal control problem
The expected cost $J:(0,T]\times\Omega\times\mathbb{A}^{\prime}\to{\mathbb{R}}$ and the value functions $V^{\prime}:(0,T]\times\Omega\to{\mathbb{R}}$ and $V:(0,T]\times\Omega\to{\mathbb{R}}$ are defined by
(4.5) $\begin{split}J(s,x,\mathbf{a})&:=\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}\Biggl{[}\int_{s}^{T}\ell(t,X^{s,x},\bar{\alpha}(t))\,{\mathrm{d}}t+h(X^{s,x})\Biggr{]},\\ V^{\prime}(s,x)&:=\inf_{\mathbf{a}\in\mathbb{A}^{\prime}}J(s,x,\mathbf{a}),\\ V(s,x)&:=\inf_{\mathbf{a}\in\mathbb{A}}J(s,x,\mathbf{a}).\end{split}$
Here, $\bar{\alpha}:{\mathbb{R}}_{+}\times\check{\Omega}\to A\cup\{\Delta^{\prime}\}$ is given by
$\bar{\alpha}(t):=\sum_{n=0}^{\infty}{\mathbf{1}}_{(T_{n},T_{n+1}]}(t)\,\mathcal{A}_{n}(t)+{\mathbf{1}}_{\left(\cup_{n=0}^{\infty}(T_{n},T_{n+1}]\right)^{c}}(t)\,a^{\circ}$
(recall $a^{\circ}$ from the beginning of section 3.2).
The proofs of the next two theorems are in section 4.5.
Theorem 4.10.
Let $(s,x)\in(0,T]\times\Omega$. Then
(4.6) $\begin{split}V^{\prime}(s,x)&=\inf_{a\in\mathcal{A}}\Bigl{\{}\chi^{s,x,a}(T)\,h(\phi^{s,x,a})+\int_{s}^{T}\Bigl{(}\chi^{s,x,a}(t)\,\ell^{s,x,a}(t)\\ &\qquad+\int_{{\mathbb{R}}^{d}}\lambda^{s,x,a}(t)\chi^{s,x,a}(t)V^{\prime}(t,\phi^{s,x,a}\otimes_{t}\,e)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr{)}\,{\mathrm{d}}t\Bigr{\}}\\ &=:[G^{\circ}V^{\prime}](s,x).\end{split}$
Theorem 4.11.
$V^{\prime}=V$.
Given Theorems 4.10 and 4.11, the proof of the following dynamic programming principle is standard (see, e.g., Proposition III.2.5 in [BardiCapuzzoDolcetta]).
Corollary 4.12.
Let $(s,x)\in(0,T]\times\Omega$ and $s_{1}\in[s,T]$ Then
(4.7) $\begin{split}V(s,x)&=\inf_{a\in\mathcal{A}}\Bigl{\{}\chi^{s,x,a}(s_{1})\,V(s_{1},\phi^{s,x,a})+\int_{s}^{s_{1}}\Bigl{(}\chi^{s,x,a}(t)\,\ell^{s,x,a}(t)\\ &\qquad+\int_{{\mathbb{R}}^{d}}\lambda^{s,x,a}(t)\chi^{s,x,a}(t)V(t,\phi^{s,x,a}\otimes_{t}\,e)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr{)}\,{\mathrm{d}}t\Bigr{\}}.\end{split}$
4.5. Proofs of Theorems 4.10 and  4.11
We will make use of an equivalent path-dependent discrete-time infinite-horizon decision model. (For corresponding work in the Markovian case, see, e.g., section 3 of [BR10PDMP], section 8.2 in [BR11book], and [RTT17PDMP].) In this section, fix $(s,x)\in(0,T]\times\Omega$.
First, we determine appropriate data for our discrete-time model. Let $\mathbf{a}\in\mathbb{A}^{\prime}$. We have
$J(s,x,\mathbf{a})$ $=\sum_{n=0}^{\infty}\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}\Bigl{[}\int_{T_{n}}^{T_{n+1}}{\mathbf{1}}_{[0,T]}(t)\,\ell(t,X^{s,x},\mathcal{A}_{n}(t))\,{\mathrm{d}}t+{\mathbf{1}}_{[T_{n},T_{n+1})}(T)\,h(X^{s,x})\Bigr{]}.$
Since, by Assumption 4.1, $h(x)=h(x(\cdot\wedge T))$ for every $x\in\Omega$, we have
${\mathbf{1}}_{[0,T]}(t_{n+1})\,h(X^{s,x}(\check{\omega}))={\mathbf{1}}_{[0,T]}(t_{n+1})\,h(X^{s,x}_{n}(\check{\omega}_{n+1}))$
for every $\check{\omega}=(\check{\omega}_{n+1},(t_{j},e_{j},\alpha_{j})_{j=n+1}^{\infty})\in\check{\Omega}$ and every $n\in{\mathbb{N}}_{0}$ (to see this, recall (4.3) for the definition of $X^{s,x}_{n}$ and Remark 4.6). Thus, for each $n\in{\mathbb{N}}_{0}$,
$\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}\Bigl{[}\int_{T_{n}}^{T_{n+1}}{\mathbf{1}}_{[0,T]}(t)\,\ell(t,X^{s,x},\mathcal{A}_{n}(t))\,{\mathrm{d}}t+{\mathbf{1}}_{[T_{n},T_{n+1})}(T)\,h(X^{s,x})\Bigr{]}$
$=\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{n+2}\Bigl{[}\int_{T_{n}\wedge T}^{T}{\mathbf{1}}_{(t,\infty]}(T_{n+1})\ell(t,X_{n}^{s,x},\mathcal{A}_{n}(t))\,{\mathrm{d}}t+{\mathbf{1}}_{[0,T]}(T_{n}){\mathbf{1}}_{(T,\infty]}(T_{n+1})\,h(X_{n}^{s,x})\Bigr{]}$
$=\int_{\check{\Omega}_{n+1}}\Big{(}\int_{t_{n}\wedge T}^{T}\chi^{t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(t)\,\ell(t,X^{s,x}_{n}(\check{\omega}_{n+1}),a_{n}(t))\,{\mathrm{d}}t$
$\qquad\qquad\qquad+{\mathbf{1}}_{[0,T]}(t_{n})\,\chi^{t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(T)\,h(X^{s,x}_{n}(\check{\omega}_{n+1}))\Big{)}\,\check{{{\mathbb{P}}}}^{s,x,\mathbf{a}}_{n+1}({\mathrm{d}}\check{\omega}_{n+1}).$
Consequently,
(4.8) $J(s,x,\mathbf{a})=\sum_{n=0}^{\infty}\check{{{\mathbb{E}}}}_{n+1}^{s,x,\mathbf{a}}\left[\vec{\ell}_{n}\right],$
where the functions $\vec{\ell}_{n}:\check{\Omega}_{n+1}\to{\mathbb{R}}_{+}$, $n\in{\mathbb{N}}_{0}$, are defined by
(4.9) $\begin{split}\vec{\ell}_{n}(\check{\omega}_{n+1})&:=\int_{t_{n}\wedge T}^{T}\chi^{t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(t)\,\ell(t,X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n}(t))\,{\mathrm{d}}t\\ &\qquad+{\mathbf{1}}_{[0,T]}(t_{n})\,\chi^{t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(T)\,h(X^{s,x}_{n}(\check{\omega}_{n+1})).\end{split}$
Here, $\check{\omega}_{n+1}=(t_{0},e_{0},\alpha_{0};\ldots;t_{n},e_{n},\alpha_{n})$.
For each $n\in{\mathbb{N}}_{0}$, define a stochastic kernel $\vec{Q}_{n}$ on (+×$({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\}$ given ((+×$(({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\})^{n+1}\times\mathcal{A}^{n+1}$ by
(+×$\int_{({\mathbb{R}}_{+}\times E)\cup\{\infty,\Delta\}}{\mathbf{1}}_{[0,\infty)}(t)\,\psi(t,e)\,\vec{Q}_{n}({\mathrm{d}}t\,{\mathrm{d}}e|(t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n})$
:={+×$:=\begin{cases}&-\int_{{\mathbb{R}}_{+}\times E}\psi(t,e)\,Q(t,{\mathrm{d}}e|t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n})\,{\mathrm{d}}F(t|t_{n},X^{s,x}_{n}(\check{\omega}_{n+1}),\alpha_{n})\\ &\qquad\qquad\text{if $\check{\omega}_{n+1}=(t_{j},e_{j},\alpha_{j})_{j=0}^{n}\in\check{\Omega}_{n+1}$,}\\ &\psi(\infty,\Delta)\qquad\text{otherwise,}\end{cases}$
for every bounded Borel-measurable function $\psi:{\mathbb{R}}_{+}\times E\cup\{(\infty,\Delta)\}\to{\mathbb{R}}$.
Remark 4.13.
Borel-measurability of $\vec{Q}_{n}$ follows from Remark 4.3 and Remark 4.6. Keep in mind that $(s,x)$ is fixed here (see the beginning of section 4.5).
Next, having determined the mentioned data, we put them into the context of section A, which is devoted to discrete-time stochastic control problems. This will allow us to use the (discrete-time) Bellman equation (A.7), which in turn will yield (4.6).
Model 4.14.
Consider a discrete-time infinite-horizon stochastic optimal control model from section A.1 with the following data:
  • State space: $S=({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\}$.
  • Control space: $U=\mathcal{A}\cup\{\Delta^{\prime}\}$.
  • Disturbance kernel $p_{k}=\vec{Q}_{k}$.
  • Cost function $g_{k}=\vec{\ell}_{k}.{\mathbf{1}}_{\check{\Omega}_{k+1}}+\infty.{\mathbf{1}}_{(S^{k+1}\times U^{k+1})\setminus\check{\Omega}_{k+1}}$ in the sense that
    $g_{k}((t_{0},e_{0}),\ldots,(t_{n},e_{n});\alpha_{0},\ldots,\alpha_{n}):=\begin{cases}\vec{\ell}_{k}((t_{j},e_{j},\alpha_{j})_{j=0}^{n})\text{ if $(t_{j},e_{j},\alpha_{j})_{j=0}^{n}\in\check{\Omega}_{n+1}$,}\\ \infty\text{ otherwise.}\end{cases}$
With the just specified data, the corresponding marginals $r_{N}(\pi,p)$ on $(S\times U)^{N}$, $N\in{\mathbb{N}}$ defined by (A.1) satisfy
$\int g_{0}\,{\mathrm{d}}r_{1}\left(\mathbf{a},\delta_{(s,x(s)}\right)$ $=\int_{S\setminus(\{0\}\times E)}\int_{U}\vec{\ell}_{0}(t_{0},e_{0},\alpha_{0})\,\mathbf{a}_{0}({\mathrm{d}}\alpha_{0})\,\delta_{(s,x(s))}({\mathrm{d}}t_{0}\,{\mathrm{d}}e_{0})$
+.$\qquad+\infty.\delta_{(s,x(s))}(\{0\}\times E)\qquad\text{(note that $s>0$)}$
$=\check{{{\mathbb{E}}}}^{s,x,\mathbf{a}}_{1}\left[\vec{\ell}_{0}\right],$
$\int g_{1}\,{\mathrm{d}}r_{2}\left(\mathbf{a},\delta_{(s,x(s))}\right)$ $=\int_{S}\int_{U}\int_{S}\int_{U}g_{1}((t_{0},e_{0}),(t_{1},e_{1});\alpha_{0},\alpha_{1})$
$\qquad\mathbf{a}_{1}({\mathrm{d}}\alpha_{1}|(t_{0},e_{0}),(t_{1},e_{1});\alpha_{0})\,\vec{Q}_{0}({\mathrm{d}}t_{1}\,{\mathrm{d}}e_{1}|(t_{0},e_{0}),\alpha_{0})$
$\qquad\mathbf{a}_{0}({\mathrm{d}}\alpha_{0})\,\delta_{(s,x(s))}({\mathrm{d}}t_{0}\,{\mathrm{d}}e_{0})$
$=\check{{{\mathbb{E}}}}^{s,\mathbf{a}}_{2}\left[\vec{\ell}_{1}\right]\qquad\text{because $\left[r_{2}\left(\mathbf{a},\delta_{(s,x(s))}\right)\right](\check{\Omega}_{2})=1$,}$
[…]
Thus, together with the definition of the optimal cost $J^{\ast}$ in (A.3), we have
(4.10) $V^{\prime}(s,x)$ $=J^{\ast}(s,x(s)).$
We need the next statement. Its proof is postponed in section B.2.
Lemma 4.15.
Fix $(s,x)\in(0,T)\times\Omega$ and $(t,e,\alpha)\in(s,T]\times E\times\mathcal{A}$. Then
$V^{\prime}(t,\phi^{s,x,\alpha}\otimes_{t}e)=J^{\ast}(1;(s,x(s)),(t,e);\alpha),$
where $J^{\ast}(1;\cdot)$ is defined by (A.6) with the data of Model 4.14.
By Theorem A.3 (below the first $\Delta^{\prime}$ can be omitted because we have $\alpha_{0}=\Delta^{\prime}$ if and only if $t_{0}=\infty$ but $t_{0}=s\in(0,T]$) and by noting that the running cost vanishes if $t>T$ due to (4.9), we have
$J^{\ast}(s,x(s))=\inf_{\alpha\in\mathcal{A}\cup\{\Delta^{\prime}\}}\Bigl{\{}g_{0}((s,x(s));\alpha)$
+(+×$\qquad\qquad+\int_{({\mathbb{R}}_{+}\times E)\cup\{(\infty,\Delta)\}}J^{\ast}(1;(s,x(s)),(t,e);\alpha)\,\vec{Q}_{0}({\mathrm{d}}t\,{\mathrm{d}}e|(s,x(s));\alpha)\Bigr{\}}$
$=\inf_{\alpha\in\mathcal{A}}\Bigl{\{}g_{0}((s,x(s));\alpha)$
$\qquad\qquad+\int_{[0,T]\times E}J^{\ast}(1;(s,x(s)),(t,e);\alpha)\,\vec{Q}_{0}({\mathrm{d}}t\,{\mathrm{d}}e|(s,x(s));\alpha)\Bigr{\}}.$
Together with (4.10) and Lemma 4.15, we get (4.6). This concludes the proof of Theorem 4.10.
By Theorem A.4,
$J^{\ast}(s,x(s))=\inf_{\mathbf{a}\in\mathbb{A}}J_{\mathbf{a}}(s,x(s)),$
where $J_{\mathbf{a}}$ is defined by (A.2). Thus
$V^{\prime}(s,x)=V(s,x).$
This concludes the proof of Theorem 4.11. ∎
5. Non-local path-dependent Hamilton-Jacobi-Bellman equations
In this section, the following additional assumption is in force. This will allow us to use the results of section 7 in [BK1].
Assumption 5.1.
The control action space $A$ is a countable union of compact metrizable subsets of $A$.
We study a terminal-value problem with a non-local path-dependent partial differential equation of the form
(5.1) $\begin{split}-\partial_{t}u(t,x)-F(t,x,u(t,x),\partial_{x}u(t,x),u(t,x\otimes_{t}\cdot))&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega,\end{split}$
where $F:[0,T]\times\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{d}\times B(E)\to{\mathbb{R}}$ is defined by
$F(t,x,y,z,\psi)$ $:=\operatornamewithlimits{essinf}_{a\in A}\Bigl{\{}\ell(t,x,a)+(f(t,x,a),z)$
$\qquad\qquad+\lambda(t,x,a)\,\int_{{\mathbb{R}}^{d}}\left[\psi(e)-y\right]\,Q(t,x,a,{\mathrm{d}}e)\Bigr{\}}.$
We aim at proving the following result (see Definition 5.10 below for our notion of minimax solution):
Theorem 5.2.
The operator $G$ defined by (5.2) has a unique fixed point $\bar{\psi}$, which is the unique minimax solution to (5.1) and coincides with the value function $V$ defined by (4.5) on $(0,T]\times\Omega$.
We continue as follows: First, we introduce the mentioned operator $G$, study its properties, and the properties of its fixed point. Then, we carry out the proof of Theorem 5.2. Finally, we establish a comparison principle between minimax sub- and supersolutions to (5.1).
5.1. The operator $G$
Let $\mathrm{BLSA}([0,T]\times\Omega)$ be the space of all bounded lower semi-analytic functions from $[0,T]\times\Omega$ to ${\mathbb{R}}$.
Given $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$ and $(s,x)\in[0,T]\times\Omega$, define
(5.2) $\begin{split}&[G\psi](s,x):=\inf_{a\in\mathcal{A}}\Bigl{\{}\chi^{s,x,a}(T)\,h(\phi^{s,x,a})\\ &+\int_{s}^{T}\chi^{s,x,a}(t)\,\Bigl{[}\ell^{s,x,a}(t)+\int_{{\mathbb{R}}^{d}}\psi(t,\phi^{s,x,a}\otimes_{t}e)\,\lambda^{s,x,a}(t)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t\Bigr{\}}.\end{split}$
Remark 5.3.
By Theorem 96, 146-IV, in [DMprobPot], Lemma 4.4, Remark 4.5, Proposition 7.48 in [BS], and Proposition 7.47 in [BS],
(5.3) $G:\mathrm{BLSA}([0,T]\times\Omega)\to\mathrm{BLSA}([0,T]\times\Omega).$
Let us introduce several operators related to $G$.
Given $s_{1}$, $s_{2}\in{\mathbb{R}}$ with $0\leq s_{1}<s_{2}\leq T$, $\eta\in\mathrm{BLSA}([s_{2},T]\times\Omega)$, and $a\in\mathcal{A}$, define131313 This is possible for the same reason that (5.3) in Remark 5.3 holds.
(5.4) $\begin{split}G_{s_{1},s_{2};\eta,a}&:\mathrm{BLSA}([s_{1},s_{2}]\times\Omega)\to\mathrm{BLSA}([s_{1},s_{2}]\times\Omega),\,\psi\mapsto(G_{s_{1},s_{2};\eta,a})\psi,\\ G_{s_{1},s_{2};\eta}&:\mathrm{BLSA}([s_{1},s_{2}]\times\Omega)\to\mathrm{BLSA}([s_{1},s_{2}]\times\Omega),\,\psi\mapsto(G_{s_{1},s_{2};\eta})\psi,\end{split}$
by
$[(G_{s_{1},s_{2};\eta,a})\psi](s,x):=\chi^{s,x,a}(s_{2})\,\eta(s_{2},\phi^{s,x,a})$
$\qquad+\int_{s}^{s_{2}}\Bigl{[}\chi^{s,x,a}(t)\,\ell^{s,x,a}(t)+\int_{{\mathbb{R}}^{d}}\psi(t,\phi^{s,x,a}\otimes_{t}e)\,\lambda^{s,x,a}(t)\chi^{s,x,a}(t)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t$
and
$[(G_{s_{1},s_{2};\eta})\psi](s,x)$ $:=\inf_{\tilde{a}\in\mathcal{A}}[(G_{s_{1},s_{2};\eta,\tilde{a}})\psi](s,x)$
for every $(s,x)\in[s_{1},s_{2}]\times\Omega$.
Theorem 5.4.
The following holds:
(i) The operator $G$ defined by (5.2) has a unique fixed point $\bar{\psi}$ in $\mathrm{BLSA}([0,T]\times\Omega)$, which coincides with $V$ on $(0,T]\times\Omega$.
(ii) If $0=r_{0}<r_{1}<\ldots<r_{N}=T$ with $r_{k+1}-r_{k}<\frac{1}{2}$ for all $k$ and if $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$, then
$(G^{n}_{r_{k-1},r_{k};\bar{\psi}})(\psi|_{[r_{k-1},r_{k}]\times\Omega})\to\bar{\psi}|_{[r_{k-1},r_{k}]\times\Omega}$
uniformly as $n\to\infty$, where the operators $G_{r_{k-1},r_{k};\bar{\psi}}$ are defined by (5.4).
Proof.
(i) Step 1 (construction of $\bar{\psi}$). Consider a partition $0=r_{0}<r_{1}<\ldots<r_{N}=T$ of $[0,T]$ with
$r_{k}-r_{k-1}<\frac{1}{2}\quad\textup{for all}\,\,k.$
Then $G_{r_{N-1},r_{N};h}$ is a contraction, say there is a $\kappa\in(0,1)$ such that, for all $\psi_{1}$, $\psi_{2}\in\mathrm{BLSA}([r_{N_{1}},r_{N}]\times\Omega)$, we have141414 If $\alpha\in\mathcal{A}$, then, for some probability measure ${{\mathbb{P}}}$, we have $[G_{r_{N-1},r_{N};h,\alpha}\,\psi_{1}](s,x)-[G_{r_{N-1},r_{N};h,\alpha}\,\psi_{2}](s,x)\leq\int_{r_{N-1}}^{r_{N}}{{\mathbb{E}}}^{{\mathbb{P}}}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}\,{\mathrm{d}}t\leq\frac{1}{2}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}.$ If $(s,x)\in[r_{N-1},r_{N}]\times\Omega$ and ${\varepsilon}>0$, then there is an $\alpha^{\varepsilon}\in\mathcal{A}$ with $[G_{r_{N-1}.r_{N};h}\,\psi_{1}](s,x)-[G_{r_{N-1}.r_{N};h}\,\psi_{2}](s,x)$ $\qquad\leq[G_{r_{N-1}.r_{N};h}\,\psi_{1}](s,x)-[G_{r_{N-1}.r_{N};h,\alpha^{\varepsilon}}\,\psi_{2}](s,x)+{\varepsilon}$ $\qquad\leq[G_{r_{N-1}.r_{N};h,\alpha^{\varepsilon}}\,\psi_{1}](s,x)-[G_{r_{N-1}.r_{N};h,\alpha^{\varepsilon}}\,\psi_{2}](s,x)+{\varepsilon}\leq\frac{1}{2}\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty}+{\varepsilon}.$
$\lVert{G_{r_{N-1},r_{N};h}\psi_{1}-G_{r_{N-1},r_{N};h}\psi_{2}}\rVert_{\infty}\leq\kappa\lVert{\psi_{1}-\psi_{2}}\rVert_{\infty},$
and thus has a unique fixed point $\bar{\psi}_{N}$. Together with (4.6), we obtain [^15] Note at this point, we have not established yet that $V^{\prime}$ is lower semi-analytic.
$\lVert{V^{\prime}|_{[r_{N-1},r_{N}]\times\Omega}-\bar{\psi}_{N}}\rVert_{\infty}$ $=\lVert{G^{\circ}V^{\prime}|_{[r_{N-1},r_{N}]\times\Omega}-G_{r_{N-1},r_{N};h}\bar{\psi}_{N}}\rVert_{\infty}$
$\leq\kappa\lVert{V^{\prime}|_{[r_{N-1},r_{N}]\times\Omega}-\bar{\psi}_{N}}\rVert_{\infty},$
i.e.,
$\bar{\psi}_{N}=V^{\prime}|_{[r_{N-1},r_{N}]\times\Omega}$
and thus $V^{\prime}$ is lower semi-analytic on $[r_{N-1},r_{N}]\times\Omega$.
Similarly, $G_{r_{N-2},r_{N-1};\bar{\psi}_{N}}$ is a contraction and thus has a unique fixed point $\bar{\psi}_{N-1}$, which coincides with $V^{\prime}$ on $[r_{N-2},r_{N-1}]\times\Omega$ thanks to Corollary 4.12.
Iterating this process yields a finite sequence $(\bar{\psi})_{k=1}^{N}$ with the properties that each $\bar{\psi}_{k}$ is the unique fixed point of $G_{r_{k-1},r_{k};\bar{\psi}_{k+1}}$, where we set
$\bar{\psi}_{N+1}(s,x):=h(x),$
and that $\bar{\psi}_{k}$ coincides with $V^{\prime}$ on $[r_{k-1},r_{k}]\times\Omega$ if $k>2$ and on $(0,r_{1}]\times\Omega$ if $k=1$.
Next, define $\bar{\psi}:[0,T]\times\Omega\to{\mathbb{R}}$ by
$\bar{\psi}(s,x):=\begin{cases}\bar{\psi}_{k}(s,x)&\text{ if $s\in[r_{k-1},r_{k})$ for some $k\in\{1,\ldots,N\}$,}\\ h(x)&\text{ if $s=T$.}\end{cases}$
Together with Theorem 4.11, we obtain $\bar{\psi}=V^{\prime}=V$ on $(0,T]\times\Omega$.
Step 2 ($\bar{\psi}$ is a fixed point of $G$). Consider the function $\bar{v}:[0,T]\times\Omega\to{\mathbb{R}}$ defined by
$\bar{v}(s,x):=[G\bar{\psi}](s,x).$
Then $\bar{v}=\bar{\psi}_{N}$ on $[r_{N-1},r_{N}]\times\Omega$.
Now, let $(s,x)\in[r_{N-2},r_{N-1})\times\Omega$. Then, by the dynamic programming principle (cf. Proposition III.2.5 in [BardiCapuzzoDolcetta]) and the fixed point property of $\bar{\psi}_{N-1}$, we have
$\bar{v}(s,x)=[(G_{r_{N-2},r_{N-1};\bar{v}})\bar{\psi}_{N-1}](s,x)=\bar{\psi}_{N-1}(s,x).$
Iterating this procedure yields $\bar{v}=\bar{\psi}$, i.e., $\bar{\psi}$ is a fixed point of $G$161616 Note this argument does not depend on properties of $V^{\prime}$ such as (4.6), which to establish was relatively technical. A simpler argument would be to just note that $V$ is lower semi-analytic by Step 1 and that then, together with (4.6), we can write $V=GV$.
Step 3 ($\bar{\psi}$ is the unique fixed point of $G$). Let $\bar{\psi}^{\prime}$ be a fixed point of $G$. On $[r_{N-1},r_{N}]\times\Omega$, we have
$\bar{\psi}^{\prime}=G\bar{\psi}^{\prime}=(G_{r_{N-1},r_{N};h})\bar{\psi}^{\prime}.$
Thus $\bar{\psi}^{\prime}$ coincides on $[r_{N-1},r_{N}]\times\Omega$ with $\bar{\psi}_{N}$, the unique fixed point of $G_{r_{N-1},r_{N};h}$.
Next, let $(s,x)\in[r_{N-2},r_{N-1}]\times\Omega$. Then, by the dynamic programming principle (cf. Proposition III.2.5 in [BardiCapuzzoDolcetta]),
$\bar{\psi}^{\prime}(s,x)$ $=[(G_{0,T;h})\bar{\psi}^{\prime}](s,x)=[(G_{r_{N-2},r_{N-1};\bar{\psi}^{\prime}})\bar{\psi}^{\prime}](s,x)=[(G_{r_{N-2},r_{N-1};\bar{\psi}_{N}})\bar{\psi}^{\prime}](s,x).$
Hence, $\bar{\psi}^{\prime}$ coincides on $[r_{N-2},r_{N-1}]\times\Omega$ with $\bar{\psi}_{N-1}$, the unique fixed point of $G_{r_{N-2},r_{N-1};\bar{\psi}_{N}}$. Iterating this process yields $\bar{\psi}^{\prime}=\bar{\psi}$.
(ii) The convergence results follow from Step 1 of part (i) of this proof. ∎
Remark 5.5.
The fixed point $\bar{\psi}$ satisfies the dynamic programming principle, i.e., if $0\leq s\leq s_{1}\leq T$ and $x\in\Omega$, then
$\bar{\psi}(s,x)=\inf_{a\in\mathcal{A}}\Bigl{\{}\int_{s}^{s_{1}}\chi^{s,x,a}(t)\,\Bigl{[}\ell^{s,x,a}(t)$
$+\int_{{\mathbb{R}}^{d}}\bar{\psi}(t,\phi^{s,x,a}\otimes_{t}e)\,\lambda^{s,x,a}(t)\,Q^{s,x,a}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t+\chi^{s,x,a}(s_{1})\bar{\psi}(s_{1},\phi^{s,x,a})\Bigr{\}}.$
Next, we show that the fixed point $\bar{\psi}$ is sufficiently regular. To simplify notation, define $\lVert{\cdot}\rVert_{s}$, $s\in[0,T]$, on $\Omega$ by
$\lVert{x}\rVert_{s}:=\sup_{0\leq t\leq s}\left|{x(t)}\right|.$
Also put
${\color[rgb]{0,0,0}\check{L}:={\mathrm{e}}^{L_{f}T}\max\{L_{f},L_{f}\,C_{f},{\color[rgb]{0,0,0}L_{f}\lVert{\bar{\psi}}\rVert_{\infty},}\,C_{f},C_{\lambda}\,L_{Q},C_{\lambda},1\}}.$
Lemma 5.6.
Let $c$, $c^{\prime}\in{\mathbb{R}}_{+}$, $0\leq s_{1}<s_{2}\leq T$, and $\psi$, $\eta\in\mathrm{BLSA}([0,T]\times\Omega)$. Let
(5.5) ${\color[rgb]{0,0,0}\left|{\psi(s,x)-\psi(s,\tilde{x})}\right|\leq c\lVert{x-\tilde{x}}\rVert_{s}}\text{ and }\left|{\eta(s,x)-\eta(s,\tilde{x})}\right|\leq c^{\prime}\lVert{x-\tilde{x}}\rVert_{s}$
for all $s\in[s_{1},s_{2}]$, $x$, $\tilde{x}\in\Omega$. Let $\lVert{\eta}\rVert_{\infty}\leq\lVert{\bar{\psi}}\rVert_{\infty}$. Recall $G_{s_{1},s_{2};\eta}$ from (5.4).
Then
(5.6) $\begin{split}&\left|{{\color[rgb]{0,0,0}[(G_{s_{1},s_{2};\eta})\psi](s,x)-[(G_{s_{1},s_{2};\eta})\psi](s,\tilde{x})}}\right|\\ &\qquad\qquad\leq{\color[rgb]{0,0,0}\left[c^{\prime}{\mathrm{e}}^{L_{f}T}+6\check{L}(s_{2}-s)(1+c)(1+\lVert{\psi}\rVert_{\infty})\right]\lVert{x-\tilde{x}}\rVert_{s}}\end{split}$
for all $(s,x,\tilde{x})\in[s_{1},s_{2}]\times\Omega\times\Omega$.
Proof.
See section B.3. ∎
Lemma 5.7.
There is a constant $L_{\bar{\psi}}\in{\mathbb{R}}_{+}$ such that, for every $s\in[0,T]$, $x$, $\tilde{x}\in\Omega$,
(5.7) ${\color[rgb]{0,0,0}\left|{\bar{\psi}(s,x)-\bar{\psi}(s,\tilde{x})}\right|\leq L_{\bar{\psi}}\lVert{x-\tilde{x}}\rVert_{s}.}$
Proof.
See section B.4. ∎
5.2. Proof of Theorem 5.2
Our approach can be considered as a path-dependent counterpart of [DavisFarid]. First, given $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$, consider the terminal-value problem
($\text{{TVP}}_{\psi}$) $\begin{split}-\partial_{t}u(t,x)-F_{\psi}(t,x,u(t,x),\partial_{x}u(t,x))&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega,\end{split}$
where
$F_{\psi}(t,x,y,z)$ $:={\color[rgb]{0,0,0}\operatornamewithlimits{essinf}}_{a\in A}\Bigl{\{}\ell(t,x,a)+(f(t,x,a),z)-\lambda(t,x,a)y$
$\qquad\qquad+\int_{{\mathbb{R}}^{d}}\psi(t,x\otimes_{t}e)\lambda(t,x,a)\,Q(t,x,a,{\mathrm{d}}e)\Bigr{\}}.$
To define our notion of solution, we need the following path spaces. Given $(s_{0},x_{0})\in[0,T)\times\Omega$ and $z\in{\mathbb{R}}^{d}$ in addition to $\psi$, define
$\mathcal{X}(s_{0},x_{0})$ :=$:=\Big{\{}x\in\Omega:\,x=x_{0}\text{ on $[0,t_{0}]$, $x|_{[t_{0},T]}$ is absolutely continuous with}$
$\qquad\qquad\left|{x^{\prime}(t)}\right|\leq C_{f}(1+\sup_{s\leq t}\left|{x(s)}\right|)\text{ a.e.~{}on $(t_{0},T)$}\Big{\}}$
and
$\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z)$ :=$:=\Big{\{}(x,y)\in\mathcal{X}(s_{0},x_{0})\times C([t_{0},T]):$
$\qquad\qquad y(t)=y_{0}+\int_{s_{0}}^{t}(x^{\prime}(s),z)-F_{\psi}(s,x,y(s),z)\,{\mathrm{d}}s\text{ on $[t_{0},T]$}\Big{\}}.$
Recall from section 3.1 that $[0,T]\times\Omega$ is considered to be equipped with the pseudo-metric defined in (3.1). Note that a function $u:[0,T]\times\Omega\to{\mathbb{R}}$ that is semicontinuous (with respect to that pseudo-metric) is non-anticipating, i.e., $u(t,x)=u(t,x_{\cdot\wedge t})$ for each $(t,x)\in[0,T]\times\Omega$.
Definition 5.8.
Let $\psi\in\mathrm{BLSA}([0,T]\times\Omega)$ and $u:[0,T]\times\Omega\to{\mathbb{R}}$.
(i) $u$ is a minimax supersolution of ($\text{{TVP}}_{\psi}$) if $u\in\mathrm{LSC}([0,T]\times\Omega)$, if $u(T,\cdot)\geq h$ on $\Omega$, and if, for every $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, and every $y_{0}\geq u(s_{0},x_{0})$, there exists an $(x,y)\in\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z)$ such that $y(t)\geq u(t,x)$ on $[s_{0},T]$.
(ii) $u$ is a minimax subsolution of ($\text{{TVP}}_{\psi}$) if $u\in\mathrm{USC}([0,T]\times\Omega)$, if $u(T,\cdot)\leq h$ on $\Omega$, and if, for every $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, and every $y_{0}\leq u(s_{0},x_{0})$, there exists an $(x,y)\in\mathcal{Y}_{\psi}(s_{0},x_{0},y_{0},z)$ such that $y(t)\leq u(t,x)$ on $[s_{0},T]$.
Theorem 5.9.
Let $\psi_{0}=\psi_{0}(s,x)\in\mathrm{BLSA}([0,T]\times\Omega)$ be Lipschitz in $x$.
(a) Then $G^{2}\psi_{0}$ is the unique minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=G\psi_{0}$.
(b) If, in addition, $\psi_{0}$ is continuous, then $G\psi_{0}$ is the unique minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=\psi_{0}$.
Proof.
(a) First, note that $G\psi_{0}$ is continuous (to see this, one can proceed as in the proof of Theorem 7.4 in [BK18JFA]). Thus, together with Theorem 96, 146-IV, in [DMprobPot], we can deduce that the maps
(5.1) $t\mapsto[G\psi_{0}](t,x\otimes_{t}e)],\,[0,T]\to{\mathbb{R}},\,(x,e)\in\Omega\times E,$
are Borel-measurable. Next, define $\ell^{[G\psi_{0}]}:[0,T]\times\Omega\times A\to{\mathbb{R}}_{+}$ by
$\ell^{[G\psi_{0}]}(t,x,a):=\ell(t,x,a)+\int_{{\mathbb{R}}^{d}}[G\psi_{0}](t,x\otimes_{t}e)\lambda(t,x,a)\,Q(t,x,a,{\mathrm{d}}e).$
By Assumptions 4.1 and 4.2, the Borel-measurability of the maps in (5.1), and the Lipschitz continuity of $\psi_{0}$ in $x$, Theorem 7.7 in [BK1] with $\ell^{[G\psi_{0}]}$ in place of $\ell$ yields part (a) of the theorem.
(b) It suffices to follow the lines of part (a) of this proof but with $G\psi_{0}$ replaced by $\psi_{0}$. This concludes the proof. ∎
Now, consider the terminal value problem
(TVP) $\begin{split}-\partial_{t}u(t,x)-F_{u}(t,x,u(t,x),\partial_{x}u(t,x))&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega.\end{split}$
Definition 5.10.
Let $u:[0,T]\times\Omega\to{\mathbb{R}}$.
(i) $u$ is a minimax supersolution of (TVP) if $u\in\mathrm{BLSA}([0,T]\times\Omega)$, $u=u(s,x)$ is Lipschitz in $x$, and $u$ is a minimax supersolution of ($\text{{TVP}}_{\psi}$) with $\psi=u$.
(ii) $u$ is a minimax subsolution of (TVP) if $u\in\mathrm{BLSA}([0,T]\times\Omega)$, $u=u(s,x)$ is Lipschitz in $x$, and $u$ is a minimax subsolution of ($\text{{TVP}}_{\psi}$) with $\psi=u$.
(iii) $u$ is a minimax solution of (TVP) if it is both a minimax supersolution and a minimax subsolution of (TVP).
Recall the fixed point $\bar{\psi}$ of the operator $G$ from Theorem 5.4.
Theorem 5.11.
$\bar{\psi}$ is the unique minimax solution of (TVP).
Proof.
(i) Existence: By Lemma 5.7, $\bar{\psi}=\bar{\psi}(s,x)$ is Lipschitz continuous in $x$. By Theorem 5.4, $\bar{\psi}\in\mathrm{BLSA}([0,T]\times\Omega)$ and
$\bar{\psi}=G\bar{\psi}=G^{2}\bar{\psi}.$
Thus, by Theorem 5.9 (a), $\bar{\psi}$ is a minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=\bar{\psi}$.
(ii) Uniqueness: Let $u$ be a minimax solution of (TVP). Then $u$ is also a minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=u$. By Theorem 5.9 (b), $Gu$ is the unique minimax solution of ($\text{{TVP}}_{\psi}$) with $\psi=u$. Moreover, by Theorem 7.7 in [BK1],
$u=Gu.$
This yields $u=\bar{\psi}$ according to Theorem 5.4. ∎
Proof of Theorem 5.2. It remains to note that Theorem 5.2 is an immediate consequence of Theorems 5.4 and 5.11. ∎
5.3. The comparison principle
Recall our notions of solution in Definition 5.10.
Theorem 5.12.
Let $u_{0}$ be a minimax sub- and $v_{0}$ be a minimax supersolution of (TVP). Then $u_{0}\leq v_{0}$.
Proof.
We adapt the idea of the proof of Theorem 5.3 in [ColaneriEtAl20SPA] to our context. Consider a partition $0=r_{0}<r_{1}<\ldots<r_{N}=T$ as in Theorem 5.4 (ii). Also recall the operators defined in (5.4).
Step 1. First, define functions
$u_{N,n}:[r_{N-1},r_{N}]\times\Omega\to{\mathbb{R}},(t,x)\mapsto u_{N,n}(t,x):=[G^{n}_{r_{N-1},r_{N};h}\,u_{0}](t,x),$
$v_{N,n}:[r_{N-1},r_{N}]\times\Omega\to{\mathbb{R}},(t,x)\mapsto v_{N,n}(t,x):=[G^{n}_{r_{N-1},r_{N};h}\,v_{0}](t,x).$
for each $n\in{\mathbb{N}}$. As $u_{0}$ and $v_{0}$ are semicontinuous, we can with the same argument as in Theorem 5.9 (b) deduce that $u_{N,1}$ (resp. $v_{N,1}$) is the unique minimax solution of ($\text{{TVP}}_{\psi}$) on $[r_{N-1},r_{N}]\times\Omega$ with $\psi=u_{0}$ (resp., $\psi=v_{0}$).
Note that $u_{0}$ (resp., $v_{0}$) is also a minimax subsolution (resp., supersolution) to ($\text{{TVP}}_{\psi}$) with $\psi=u_{0}$. Thus, by the comparison principle for ($\text{{TVP}}_{\psi}$) (Corollary  5.4 in [BK1]), $u_{0}\leq u_{N,1}$ and $v_{N,1}\leq v_{0}$ on $[r_{N-1},r_{N}]\times\Omega$.
By induction, one can similarly show that $(u_{N,n})$ is non-decreasing and that $(v_{N,n})$ is non-increasing. Moreover, by Theorem 5.4 (ii), $u_{N,n}\to\bar{\psi}$ and $v_{N,n}\to\bar{\psi}$ uniformly on $[r_{N-1},r_{N}]\times\Omega$ as $n\to\infty$. Thus $u_{0}\leq\bar{\psi}\leq v_{0}$ on $[r_{N-1},r_{N}]\times\Omega$.
Step 2. Let us next consider the interval $[r_{N-2},r_{N}]$. We can proceed as in Step 1 and arrive at the corresponding conclusion but need to replace $h$ by $\bar{\psi}$.
Step 3. Iterating the steps above yields $u_{0}\leq\bar{\psi}\leq v_{0}$ on $[0,T]\times\Omega$. ∎
Appendix A A general path-dependent discrete-time infinite-horizon decision model
We set up a discrete-time infinite-horizon decision model with data that depend on past states and past controls. Then we formulate a corresponding discrete-time infinite-horizon stationary decision model (see also chapter 10 of [BS] regarding a related procedure concerning non-stationary models, but note that our situation is slightly more involved). Finally, we establish part of the Bellman equation for the former model by using the latter model.
A.1. A model with data that depend on past states and past controls
Our notation and terminology follows largely [BS].
Consider a model with the following data (cf. Definition 8.1 on p. 188 in [BS]):
  • State space $S$, a non-empty Borel space, which also serves as disturbance space.
  • Control space $U$, a non-empty Borel space.
  • Disturbance kernel
    $p_{k}({\mathrm{d}}w|x_{0},\ldots,x_{k};u_{0},\ldots,u_{k})\quad\textup{at stage}\,\,k,\quad k\in{\mathbb{N}}_{0},$
    a Borel-measurable stochastic kernel on $S$ given $S^{k+1}\times U^{k+1}$.
  • System function
    $(x,u,w)\mapsto w,\quad S\times U\times S\to S.$
  • Discount factor $1$.
  • Cost function
    $g_{k}:S^{k+1}\times U^{k+1}\to{\mathbb{R}}_{+}\cup\{+\infty\}\quad\textup{at stage}\,\,k,\quad k\in{\mathbb{N}}_{0},$
    a lower semi-analytic function.
Remark A.1.
Maybe more suggestive (and more in line with [BS]), we can write that our system function maps the $k$-th disturbance $w_{k}$ to the $(k+1)$-st state $x_{k+1}$.
Our goal is to minimize a certain expected cost over the space $\Pi^{\prime}$ of all policies $\pi=(\mu_{0},\mu_{1},\ldots)$, where each
$\mu_{k}({\mathrm{d}}u_{k}|x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1}),\quad k\in{\mathbb{N}}_{0},$
is a universally measurable stochastic kernel on $U$ given $S^{k+1}\times U^{k}$ (note that $\mu_{0}$ is a kernel on $U$ given $S$).
Remark A.2.
We use randomized policies for two reasons. First, by doing so, we operate immediately in the setting of [BS], whose results we need to use. Second, the notation for non-randomized policies would become very cumbersome in our model. E.g., (realized) policies would then be of the form $\mu_{0}(x_{0})$, $\mu_{1}(x_{0},x_{1};\mu_{0}(x_{0}))$, $\mu_{2}(x_{0},x_{1},x_{2};\mu_{0}(x_{0}),\mu_{1}(x_{0},x_{1};\mu_{0}(x_{0})))$, and so on. The randomized counterparts are of the form $\mu_{0}({\mathrm{d}}u_{0}|x_{0})$, $\mu_{1}({\mathrm{d}}u_{1}|x_{0},x_{1};u_{0})\,\mu_{0}({\mathrm{d}}u_{0}|x_{0})$, $\ldots$,
$\mu_{n+1}({\mathrm{d}}u_{n+1}|x_{0},\ldots,x_{n+1};u_{0},\ldots,u_{n})\cdots\mu_{0}({\mathrm{d}}u_{0}|x_{0}).$
We consider the notation for the latter (realized) policies to be slightly less heavy than the notation for the former ones.
Given a policy $\pi=(\mu_{0},\mu_{1},\ldots)\in\Pi^{\prime}$ and a probability measure $p$ on $S$, define a probability measure $r(\pi,p)$ on $(S\times U)^{{\mathbb{N}}_{0}}$ via marginals $r_{N}(\pi,p)$ on $(S\times U)^{N}$, $N\in{\mathbb{N}}$, that are recursively defined by
(A.1) $\begin{split}\int h_{0}\,{\mathrm{d}}r_{1}(\pi,p)&:=\int_{S}\int_{U}h_{0}(x_{0};u_{0})\,\mu_{0}({\mathrm{d}}u_{0}|x_{0})\,p({\mathrm{d}}x_{0}),\\ \int h_{1}\,{\mathrm{d}}r_{2}(\pi,p)&:=\int_{S}\int_{U}\int_{S}\int_{U}h_{1}(x_{0},x_{1};u_{0},u_{1})\\ &\qquad\qquad\mu_{1}({\mathrm{d}}u_{1}|x_{0},x_{1};u_{0})\,p_{0}({\mathrm{d}}x_{1}|x_{0},u_{0})\\ &\qquad\qquad\mu_{0}({\mathrm{d}}u_{0}|x_{0})\,p({\mathrm{d}}x_{0}),\\ \int h_{N}\,{\mathrm{d}}r_{N+1}(\pi,p)&:=\int\Bigl{[}\int_{S}\int_{U}h_{N}(x_{0},\ldots,x_{N};u_{0},\ldots,u_{N})\\ &\qquad\qquad\mu_{N}({\mathrm{d}}u_{N}|x_{0},\ldots,x_{N};u_{0},\ldots,u_{N-1})\\ &\qquad\qquad p_{N-1}({\mathrm{d}}x_{N}|x_{0},\ldots,x_{N-1};u_{0},\ldots,u_{N-1})\Bigr{]}\\ &\qquad r_{N}(\pi,p)({\mathrm{d}}x_{0}\,{\mathrm{d}}u_{0}\,\ldots\,{\mathrm{d}}x_{N-1}\,{\mathrm{d}}u_{N-1}),\,N\in{\mathbb{N}},\end{split}$
for each bounded universally measurable function $h_{k}:S^{k+1}\times U^{k+1}\to{\mathbb{R}}$, $k\in{\mathbb{N}}_{0}$.
Given a policy $\pi\in\Pi^{\prime}$ and a state $x\in S$, define the cost
(A.2) $\begin{split}J_{\pi}(x)&:=\int\sum_{k=0}^{\infty}g_{k}\,{\mathrm{d}}r(\pi,\delta_{x})\\ &=\sum_{k=0}^{\infty}\int g_{k}(x_{0},\ldots,x_{k};u_{0},\ldots,u_{k})\,r_{k+1}(\pi,\delta_{x})({\mathrm{d}}x_{0}\,{\mathrm{d}}u_{0}\,\ldots\,{\mathrm{d}}x_{k}\,{\mathrm{d}}u_{k}).\end{split}$
Given a state $x\in S$, define the optimal cost
(A.3) $J^{\ast}(x):=\inf_{\pi\in\Pi^{\prime}}J_{\pi}(x).$
We also use the spaces $\Pi^{k}$, $k\in{\mathbb{N}}_{0}$, of all policies $\pi^{k}=(\mu_{k},\mu_{k+1},\ldots)$, where each
$\mu_{j}({\mathrm{d}}u_{j}|x_{0},\ldots,x_{j};u_{0},\ldots,u_{j-1}),\quad j\in\{k,k+1,\ldots\},$
is a universally measurable stochastic kernel on $U$ given $S^{j+1}\times U^{j}$ (cf. chapter 10 of [BS]). Note that $\Pi^{\prime}=\Pi^{0}$.
Next, given $k\in{\mathbb{N}}_{0}$, a policy $\pi^{k}=(\mu_{k},\mu_{k+1},\ldots)\in\Pi^{k}$, and a probability measure $p^{k}$ on $S^{k+1}\times U^{k}$, define a probability measure $r^{k}(\pi^{k},p^{k})$ on $(S\times U)^{{\mathbb{N}}_{0}}$ via marginals $r^{k}_{N}(\pi^{k},p^{k})$ on $(S\times U)^{N}$, $N\in\{k+1,k+2,\ldots\}$, that are recursively defined by
(A.4) $\begin{split}\int h_{k}\,{\mathrm{d}}r^{k}_{k+1}(\pi^{k},p^{k})&:=\int_{S^{k+1}\times U^{k}}\int_{U}h_{k}(x_{0},\ldots,x_{k};u_{0},\ldots,u_{k})\,\\ &\qquad\qquad\mu_{k}({\mathrm{d}}u_{k}|x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1})\\ &\qquad\qquad p^{k}({\mathrm{d}}x_{0}\,\ldots\,{\mathrm{d}}x_{k}\,{\mathrm{d}}u_{0},\ldots\,{\mathrm{d}}u_{k-1}),\\ \int h_{N}\,{\mathrm{d}}r^{k}_{N+1}(\pi^{k},p^{k})&:=\int\Big{[}\int_{S}\int_{U}h_{N}(x_{0},\ldots,x_{N};u_{0},\ldots,u_{N})\\ &\qquad\qquad\mu_{N}({\mathrm{d}}u_{N}|x_{0},\ldots,x_{N};u_{0},\ldots,u_{N-1})\\ &\qquad\qquad p_{N-1}({\mathrm{d}}x_{N}|x_{0},\ldots,x_{N-1};u_{0},\ldots,u_{N-1})\Big{]}\\ &\,r^{k}_{N}(\pi^{k},p^{k})({\mathrm{d}}x_{0}\,{\mathrm{d}}u_{0}\,\ldots\,{\mathrm{d}}x_{N-1}\,{\mathrm{d}}u_{N-1})\end{split}$
for every bounded universally measurable function $h_{j}:S^{j+1}\times U^{j+1}\to{\mathbb{R}}$, $j\in\{k,k+1,\ldots\}$.
Given $k\in{\mathbb{N}}_{0}$, $\pi^{k}\in\Pi^{k}$, $x_{0}$, $\ldots$, $x_{k}\in S$, and $u_{0}$, $\ldots$, $u_{k-1}\in U$, define the cost
(A.5) $J_{\pi^{k}}(k;x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1})$ $:=\int\sum_{j=k}^{\infty}g_{j}\,{\mathrm{d}}r^{k}(\pi^{k},\delta_{(x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1})}).$
Given $k\in{\mathbb{N}}_{0}$, $x_{0}$, $\ldots$, $x_{k}\in S$, and $u_{0}$, $\ldots$, $u_{k-1}\in U$, define the optimal cost
(A.6) $J^{\ast}(k;x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1})):=\inf_{\pi^{k}\in\Pi^{k}}J_{\pi^{k}}(k;x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1}).$
Our goal is to establish the following special case of the Bellman equation (this is a counterpart of Proposition 10.1 on p. 246 in [BS]).
Theorem A.3.
Let $x\in S$. Let $J^{\ast}$ and $J^{\ast}(1;\cdot)$ be the functions defined respectively by (A.3) and (A.6) with $k=1$. Then
(A.7) $J^{\ast}(x)=\inf_{u\in U}\Bigl{\{}g_{0}(x,u)+\int_{S}J^{\ast}(1;x,x_{1};u)\,p_{0}({\mathrm{d}}x_{1}|x,u)\Bigr{\}}.$
Proof.
See section A.3. ∎
Moreover, we show that minimization over non-randomized policies is equivalent to minimization over randomized policies. We want to emphasize that we still work with the data introduced at the beginning of this section (section A.1). In particular, non-negativity of the cost functions is important. This will enable us to apply corresponding results in [BS].
Theorem A.4.
Let $x\in S$. Let $J^{\ast}$ and $J_{\pi}$ be the functions defined respectively by (A.3) and (A.2) Then
(A.8) $J^{\ast}(x)=\inf_{\pi\in\Pi}J_{\pi}(x),$
where $\Pi$ is the set of all non-randomized policies $\pi=(\mu_{0},\mu_{1},\ldots)\in\Pi^{\prime}$, i.e., for every $k\in{\mathbb{N}}_{0}$, every $x_{0}$, $\ldots$, $x_{k}\in S$, and every $u_{0}$, $\ldots$, $u_{k-1}\in U$, the measure $\mu_{k}(du_{k}|x_{0},\ldots,x_{k};u_{0},\ldots u_{k-1})$ is a Dirac measure.
Proof.
See section A.4. ∎
A.2. A corresponding stationary model
Using the data from the previous subsection, consider a model with the following data:
  • State space $\mathbf{S}:={\mathbb{N}}_{0}\times S^{{\mathbb{N}}_{0}}\times U^{{\mathbb{N}}_{0}}$. A typical state is denoted by
    $\mathbf{x}=(k;x^{0},x^{1},\ldots;u^{0},u^{1},\ldots)$
    and $\mathbf{x}$ is to be understood of that form below.
  • Control space $U$.
  • Disturbance space $S$.
  • Disturbance kernel
    $\mathbf{p}({\mathrm{d}}w|\mathbf{x},u):=p_{k}({\mathrm{d}}w|x^{0},\ldots,x^{k};u^{0},\ldots,u^{k-1},u)$
    on $\mathbf{S}$ given $\mathbf{S}\times U$.
  • System function $\mathbf{f}:\mathbf{S}\times U\times S\to\mathbf{S}$ defined by
    $\mathbf{f}(\mathbf{x},u,w):=(k+1;x^{0},\ldots,x^{k},w,x^{k+2},x^{k+3},\ldots;u^{0},\ldots,u^{k-1},u,u^{k+1},u^{k+2},\ldots).$
  • Discount factor $1$.
  • Cost function $\mathbf{g}:\mathbf{S}\times U\to{\mathbb{R}}_{+}$ defined by
    $\mathbf{g}(\mathbf{x},u):=g_{k}(x^{0},\ldots,x^{k};u^{0},\ldots,u^{k-1},u).$
The corresponding state transition kernel $\mathbf{t}$ on $\mathbf{S}$ given $\mathbf{S}\times U$ is (as on p. 189 in [BS]) defined by
(A.9) $\begin{split}\mathbf{t}(\underline{\mathbf{S}}|\mathbf{x},u)&:=\mathbf{p}(\{w\in S:\,\mathbf{f}(\mathbf{x},u,w)\in\underline{\mathbf{S}}\}|\mathbf{x},u)\\ &=\delta_{k+1}(\underline{E})\cdot\prod_{j=0}^{k}\delta_{x^{j}}(\underline{F}_{j})\cdot p_{k}(\underline{F}_{k+1}|x^{0},\ldots,x^{k};u^{0},\ldots,u^{k-1},u)\\ &\qquad\qquad\cdot\prod_{j={k+2}}^{\infty}\delta_{x^{j}}(\underline{F}_{j})\cdot\prod_{j=0}^{k-1}\delta_{u^{j}}(\underline{G}_{j})\cdot\delta_{u}(\underline{G}_{k})\cdot\prod_{j=k+1}^{\infty}\delta_{u^{j}}(\underline{G}_{j}),\end{split}$
for each $\underline{\mathbf{S}}=\underline{E}\times\prod_{j=0}^{\infty}\underline{F}_{j}\times\prod_{j=0}^{\infty}\underline{G}_{j}\in\mathcal{B}(\mathbf{S})$, $\mathbf{x}=(k;x^{0},x^{1},\ldots;u^{0},u^{1},\ldots)\in\mathbf{S}$, and $u\in U$.
The set of policies, denoted by $\vec{\Pi}^{\prime}$, consists of all $\vec{\pi}=(\vec{\mu}_{0},\vec{\mu}_{1},\ldots)$, where each
$\vec{\mu}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\,\ldots,\,\mathbf{x}_{k};\,u_{0},\,\ldots,u_{k-1}),\quad k\in{\mathbb{N}}_{0},$
is a universally measurable stochastic kernel on $U$ given $\mathbf{S}^{k+1}\times U^{k}$.
The set of all Markov policies, denoted by $\vec{\Pi}$, consists of all $\vec{\pi}=(\vec{\mu}_{0},\vec{\mu}_{1},\ldots)\in\vec{\Pi}^{\prime}$, where each $\vec{\mu}_{k}$, $k\in{\mathbb{N}}_{0}$, only depends on $\mathbf{x}_{k}$ instead of $\mathbf{x}_{0}$, $\ldots$, $\mathbf{x}_{k}$, $u_{0}$, $\ldots$, $u_{k-1}$, and in this case we write
$\vec{\mu}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{k})$
instead of $\vec{\mu}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\,\ldots,\,\mathbf{x}_{k};\,u_{0},\,\ldots,u_{k-1})$.
For each policy $\vec{\pi}=(\vec{\mu}_{0},\vec{\mu}_{1},\ldots)\in\vec{\Pi}^{\prime}$ and each probability measure $\mathbf{p}$ on $\mathbf{S}$, define a probability measure $\mathbf{r}(\vec{\pi},\mathbf{p})$ on $(\mathbf{S}\times U)^{{\mathbb{N}}_{0}}$ via marginal distributions $\mathbf{r}_{N}(\vec{\pi},\mathbf{p})$ on $(\mathbf{S}\times U)^{N}$, $N\in{\mathbb{N}}$, by
(A.10) $\begin{split}\int\mathbf{h}_{0}\,{\mathrm{d}}\mathbf{r}_{1}(\vec{\pi},\mathbf{p})&:=\int_{\mathbf{S}}\int_{U}\mathbf{h}_{0}(\mathbf{x}_{0};u_{0})\,\vec{\mu}_{0}({\mathrm{d}}u_{0}|\mathbf{x}_{0})\,\mathbf{p}({\mathrm{d}}\mathbf{x}_{0}),\\ \int\mathbf{h}_{N+1}\,{\mathrm{d}}\mathbf{r}_{N+1}(\vec{\pi},\mathbf{p})&:=\int\Biggl{[}\int_{\mathbf{S}}\int_{U}\mathbf{h}_{N}(\mathbf{x}_{0},\ldots,\mathbf{x}_{N};u_{0},\ldots,u_{N})\\ &\qquad\qquad\vec{\mu}_{N}({\mathrm{d}}u_{N}|\mathbf{x}_{0},\ldots,\mathbf{x}_{N};u_{0},\ldots,u_{N-1})\\ &\qquad\qquad\mathbf{t}({\mathrm{d}}\mathbf{x}_{N}|\mathbf{x}_{N-1},u_{N-1})\Biggr{]}\\ &\qquad\mathbf{r}_{N}(\vec{\pi},\mathbf{p})({\mathrm{d}}\mathbf{x}_{0}\,{\mathrm{d}}u_{0}\,\ldots\,{\mathrm{d}}\mathbf{x}_{N-1}\,{\mathrm{d}}u_{N-1}),\,N\in{\mathbb{N}},\end{split}$
for every bounded universally measurable function $\mathbf{h}_{k}:\mathbf{S}^{k+1}\times U^{k+1}\to{\mathbb{R}}$, $k\in{\mathbb{N}}_{0}$.
Now, we can define the cost functions $\mathbf{J}_{\vec{\pi}}:\mathbf{S}\to{\mathbb{R}}$, $\vec{\pi}\in\vec{\Pi}^{\prime}$, and the optimal cost function $\mathbf{J}:\mathbf{S}\to{\mathbb{R}}$ by
(A.11) $\begin{split}\mathbf{J}_{\vec{\pi}}(\mathbf{x})&:=\sum_{k=0}^{\infty}\int\mathbf{g}(\mathbf{x}_{k},u_{k})\,\mathbf{r}_{k+1}(\vec{\pi},\delta_{\mathbf{x}})({\mathrm{d}}\mathbf{x}_{0}\,{\mathrm{d}}u_{0}\,\ldots\,{\mathrm{d}}\mathbf{x}_{k}\,{\mathrm{d}}u_{k}),\\ \mathbf{J}^{\ast}(\mathbf{x})&:=\inf_{\vec{\pi}\in\vec{\Pi}^{\prime}}\mathbf{J}_{\vec{\pi}}(\mathbf{x}).\end{split}$
A.3. Proof of Theorem A.3.
Recall that $\mathbf{g}$ is non-negative. Thus, by Proposition 9.8 on p. 225 in [BS] and Definition 8.5 on p. 195 in [BS], we have
(A.12) $\mathbf{J}^{\ast}(\mathbf{x})=\inf_{u\in U}\Bigl{\{}\mathbf{g}(\mathbf{x},u)+\int\mathbf{J}^{\ast}(\mathbf{x}^{\prime})\,\mathbf{t}({\mathrm{d}}\mathbf{x}^{\prime}|\mathbf{x},u)\Bigr{\}}$
for every $\mathbf{x}\in\mathbf{S}$.
Now, fix $x\in S$ and some $u_{\Delta}\in U$, and let $\mathbf{x}=(0;x,x\ldots;u_{\Delta},u_{\Delta},\ldots)$. By (A.12),
$\mathbf{J}^{\ast}(\mathbf{x})=\inf_{u\in U}\Bigl{\{}g_{0}(x;u)+\int\mathbf{J}^{\ast}(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots)\,p_{0}({\mathrm{d}}w|x,u)\Bigr{\}}.$
Now, fix also $u\in U$ and $w\in S$. It remains to show that
(A.13) $\begin{split}\mathbf{J}^{\ast}(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots)&=J^{\ast}(x)\quad\text{and}\\ \mathbf{J}^{\ast}(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots)&=J^{\ast}(1;x,w;u),\end{split}$
which we will do by proving the following four claims. Note that we will use the following notational conventions for the components of elements $(\mathbf{x}_{0},\mathbf{x}_{1},\ldots)$ of $\mathbf{S}^{{\mathbb{N}}_{0}}$:
(A.14) $\mathbf{x}_{k}=(\iota_{k};x^{0}_{k},x^{1}_{k},\ldots;u^{0}_{k},u^{1}_{k},\ldots),\quad k\in{\mathbb{N}}_{0}.$
Claim 1: $\mathbf{J}^{\ast}(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots)\geq J^{\ast}(x)$.
Claim 2: $\mathbf{J}^{\ast}(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots)\geq J^{\ast}(1;x,w;u)$.
Claim 3: $\mathbf{J}^{\ast}(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots)\leq J^{\ast}(x)$.
Claim 4: $\mathbf{J}^{\ast}(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots)\leq J^{\ast}(1;x,w;u)$.
Proof of Claim 1. Fix a policy $\vec{\pi}=(\vec{\mu}_{0},\vec{\mu}_{1},\ldots)\in\vec{\Pi}^{\prime}$. Next, define a policy $\pi=(\mu_{0},\mu_{1},\ldots)\in\Pi^{\prime}$ as follows. For every $k\in{\mathbb{N}}_{0}$, put
$\mu_{k}({\mathrm{d}}u_{k}|x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1})$
$:=\vec{\mu}_{k}({\mathrm{d}}u_{k}|(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots),(1;x,x_{1},x,x,\ldots;u_{0},u_{\Delta},u_{\Delta},\ldots),$
$\qquad\qquad(2;x,x_{1},x_{2},x,x,\ldots;u_{0},u_{1},u_{\Delta},u_{\Delta},\ldots),\ldots,$
$\qquad\qquad(k;x,x_{1},\ldots,x_{k},x,x,\ldots;u_{0},\ldots,u_{k-1},u_{\Delta},u_{\Delta},\ldots);u_{0},\ldots,u_{k-1}).$
By (A.11), (A.10) (cf. also (4) on p. 191 in [BS]), (A.9), and (A.2),
$\mathbf{J}_{\vec{\pi}}(\mathbf{x})$ $=\sum_{k=0}^{\infty}\int_{\mathbf{S}}\int_{U}\int_{\mathbf{S}}\int_{U}\cdots\int_{\mathbf{S}}\int_{U}\mathbf{g}(\mathbf{x}_{k},u_{k})\,\vec{\mu}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\ldots,\mathbf{x}_{k};u_{0},\ldots,u_{k-1})$
$\qquad\mathbf{t}({\mathrm{d}}\mathbf{x}_{k}|\mathbf{x}_{k-1},u_{k-1})\cdots\vec{\mu}_{1}({\mathrm{d}}u_{1}|\mathbf{x}_{0},\mathbf{x}_{1};u_{0})\,\mathbf{t}({\mathrm{d}}\mathbf{x}_{1}|\mathbf{x}_{0},u_{0})\,\vec{\mu}_{0}({\mathrm{d}}u_{0}|\mathbf{x}_{0})\,\delta_{\mathbf{x}}({\mathrm{d}}\mathbf{x}_{0})$
$=\sum_{k=0}^{\infty}\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{k}(x,x^{1}_{1},\ldots,x^{k}_{k};u_{0},\ldots,u_{k-1},u_{k})$
$\qquad\vec{\mu}_{k}({\mathrm{d}}u_{k}|(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots),(1;x,x_{1},x,x,\ldots;u_{0},u_{\Delta},u_{\Delta},\ldots),$
$\qquad\qquad(2;x,x^{1}_{1},x^{2}_{2},x,x,\ldots;u_{0},u_{1},u_{\Delta},u_{\Delta},\ldots),\ldots,$
$\qquad\qquad(k;x,x^{1}_{1},\ldots,x^{k}_{k},x,x,\ldots;u_{0},\ldots,u_{k-1},u_{\Delta},u_{\Delta},\ldots);u_{0},\ldots,u_{k-1})$
$\qquad p_{k-1}({\mathrm{d}}x^{k}_{k}|x,x^{1}_{1},\ldots,x^{k-1}_{k-1};u_{0},\ldots,u_{k-1})\cdots$
$\qquad\vec{\mu}_{1}({\mathrm{d}}u_{1}|(0;x,x\ldots;u_{\Delta},u_{\Delta},\ldots),(1;x,x_{1}^{1},x,x,\ldots;u_{0},u_{\Delta},u_{\Delta},\ldots);u_{0})$
$\qquad p_{0}({\mathrm{d}}x_{1}^{1}|x,u_{0})\,\vec{\mu}_{0}({\mathrm{d}}u_{0}|(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots))$
$=J_{\pi}(x).$
This yields Claim 1.
Proof of Claim 2. Fix a policy $\vec{\pi}=(\vec{\mu}_{0},\vec{\mu}_{1},\ldots)\in\vec{\Pi}^{\prime}$ and define a policy $\pi^{1}=(\nu_{1},\nu_{2},\ldots)\in\Pi^{1}$ as follows. For every $k\in{\mathbb{N}}_{0}$, put
$\nu_{k+1}({\mathrm{d}}u_{k+1}|x_{0},\ldots,x_{k+1};u_{0},\ldots,u_{k})$
$:=\vec{\mu}_{k}({\mathrm{d}}u_{k+1}|(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots),(2;x,w,x_{2},x,x,\ldots;u,u_{1},u_{\Delta},u_{\Delta},\ldots),$
$\qquad(3;x,w,x_{2},x_{3},x,x,\ldots;u,u_{1},u_{2},u_{\Delta},u_{\Delta},\ldots),\ldots,$
$\qquad(k+1;x,w,x_{2},\ldots,x_{k+1},x,x,\ldots;u,u_{1},\ldots,u_{k},u_{\Delta},u_{\Delta},\ldots);u_{1},\ldots,u_{k}).$
such that
$\nu_{k+1}({\mathrm{d}}u_{k}|x,w,x_{2},\ldots,x_{k+1};u,u_{0},\ldots,u_{k-1})$
$=\vec{\mu}_{k}({\mathrm{d}}u_{k}|(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots),(2;x,w,x_{2},x,x,\ldots;u,u_{0},u_{\Delta},u_{\Delta},\ldots),$
$\qquad(3;x,w,x_{2},x_{3},x,x,\ldots;u,u_{0},u_{1},u_{\Delta},u_{\Delta},\ldots),\ldots,$
$\qquad(k+1;x,w,x_{2},\ldots,x_{k+1},x,x,\ldots;u,u_{0},\ldots,u_{k-1},u_{\Delta},u_{\Delta},\ldots);u_{0},\ldots,u_{k-1}).$
Hence, by (A.11), (A.10), (A.9), and (A.5),
$\mathbf{J}_{\vec{\pi}}(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots)$
$=\sum_{k=0}^{\infty}\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{k+1}(x,w,x^{2}_{1},\ldots,x^{k+1}_{k};u,u_{0},u_{1},\ldots,u_{k-2},u_{k-1},u_{k})$
$\quad\vec{\mu}_{k}({\mathrm{d}}u_{k}|(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots),(2;x,w,x^{2}_{1},x,x,\ldots;u,u_{0},u_{\Delta},u_{\Delta},\ldots),$
$\quad\quad(3;x,w,x,x^{2}_{1},x^{3}_{2},x,x,\ldots;u,u_{0},u_{1},u_{\Delta},u_{\Delta},\ldots),\ldots,$
$\quad\quad(k+1;x,w,x,x^{2}_{1},\ldots,x^{k+1}_{k},x,x,\ldots;u,u_{0},\ldots,u_{k-1},u_{\Delta},u_{\Delta},\ldots);u_{0},\ldots,u_{k-1})$
$\quad p_{k}({\mathrm{d}}x^{k+1}_{k}|x,w,x^{2}_{1},\ldots,x^{k}_{k-1};u,u_{0},\ldots,u_{k-1})\cdots$
$\quad\vec{\mu}_{1}({\mathrm{d}}u_{1}|(1;x,w,x,x\ldots;u,u_{\Delta},u_{\Delta},\ldots),(2;x,w,x_{1}^{2},x,x,\ldots;u,u_{0},u_{\Delta},u_{\Delta},\ldots);u_{0})$
$\quad p_{1}({\mathrm{d}}x_{1}^{2}|x,w;u,u_{0})\,\vec{\mu}_{0}({\mathrm{d}}u_{0}|(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots)),$
where it might be also helpful to note that, by (A.9) and (A.14), we have
$\mathbf{t}({\mathrm{d}}\mathbf{x}_{1}|(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots);u_{0})=\delta_{2}({\mathrm{d}}\iota_{1})\,\delta_{x}({\mathrm{d}}x^{0}_{1})\,\delta_{w}({\mathrm{d}}x^{1}_{1})$
$\qquad\cdot p_{1}({\mathrm{d}}x^{2}_{1}|x,w;u,u_{0})\,\prod_{j=3}^{\infty}\delta_{x}({\mathrm{d}}x^{j}_{1})\,\delta_{u}({\mathrm{d}}u^{0}_{1})\,\delta_{u_{0}}({\mathrm{d}}u^{1}_{1})\,\prod_{j=2}^{\infty}\delta_{u_{\Delta}}({\mathrm{d}}u^{j}_{1}),\text{ etc.}$.
Next, changing the lower limit of summation and then relabeling our integration variables, we obtain
$\mathbf{J}_{\vec{\pi}}(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots)$
$=\sum_{k=1}^{\infty}\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{k}(x,w,x^{2},\ldots,x_{k-1}^{k};u,u_{0},u_{k-1})$
$\qquad\qquad\nu_{k}({\mathrm{d}}u_{k-1}|x,w,x^{2}_{1},\ldots,x^{k}_{k-1};u,u_{0},\ldots,u_{k-2})$
$\qquad\qquad p_{k-1}({\mathrm{d}}x^{k}_{k-1}|x,w,x^{2}_{1},\ldots,x^{k-1}_{k-2};u,u_{0},\ldots,u_{k-2})\cdots$
$\qquad\qquad\nu_{2}({\mathrm{d}}u_{1}|x,w,x^{2}_{1};u,u_{0})\,p_{1}({\mathrm{d}}x^{2}_{1}|x,w;u,u_{0})\,\nu_{1}({\mathrm{d}}u_{0}|x,w;u)$
$=\sum_{k=1}^{\infty}\int_{S}\int_{U}\int_{S}\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{k}(x_{0},\ldots,x_{k};u_{0},\ldots,u_{k})$
$\qquad\qquad\nu_{k}({\mathrm{d}}u_{k}|x_{0},\ldots,x_{k};u_{0},\ldots,u_{k-1})\,p_{k-1}({\mathrm{d}}x_{k}|x_{0},\ldots,x_{k-1};u_{0},\ldots,u_{k-1})\cdots$
$\qquad\qquad\nu_{2}({\mathrm{d}}u_{1}|x_{0},x_{1},x_{2};u_{0},u_{1})\,p_{1}({\mathrm{d}}x_{2}|x_{0},x_{1};u_{0},u_{1})\,\nu_{1}({\mathrm{d}}u_{1}|x_{0},x_{1};u_{0})$
$\qquad\qquad\delta_{w}({\mathrm{d}}x_{1})\,\delta_{u}({\mathrm{d}}u_{0})\,\delta_{x}({\mathrm{d}}x_{0})$
$=J_{\pi^{1}}(1;x,w;u).$
This yields Claim 2.
Proof of Claim 3. Fix a policy $\pi=(\mu_{0},\mu_{1},\ldots)\in\Pi^{\prime}$. Define a policy $\vec{\pi}=(\vec{\pi}_{0},\vec{\pi}_{1},\ldots)\in\vec{\Pi}^{\prime}$ by
$\vec{\mu}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\ldots,\mathbf{x}_{k};u_{0},\ldots,u_{k-1}):=\mu_{k}({\mathrm{d}}u_{k}|x_{0}^{0},\ldots,x_{k}^{k};u_{0},\ldots,u_{k-1}),\quad k\in{\mathbb{N}}_{0}.$
Then, similarly as in the proof of Claim 1, we have
$J_{\pi}(x)$ $=\sum_{k=0}^{\infty}\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{k}(x,x^{1}_{1},\ldots,x^{k}_{k};u_{0},\ldots,u_{k-1},u_{k})$
$\qquad\mu_{k}({\mathrm{d}}u_{k}|x,x^{1}_{1},\ldots,x^{k}_{k};u_{0},\ldots,u_{k-1})$
$\qquad p_{k-1}({\mathrm{d}}x^{k}_{k}|x,x^{1}_{1},\ldots,x^{k-1},x_{k-1};u_{0},\ldots,u_{k-1})\cdots$
$\qquad\mu_{1}({\mathrm{d}}u_{1}|x,x^{1}_{1};u_{0})\,p_{0}({\mathrm{d}}x^{1}_{1}|x;u_{0})\,\mu_{0}({\mathrm{d}}u_{0}|x)$
$=\mathbf{J}_{\vec{\pi}}(\mathbf{x}).$
This yields Claim 3.
Proof of Claim 4. Fix a policy $\pi^{1}=(\mu_{1},\mu_{2},\ldots)\in\Pi^{1}$. Define a policy $\vec{\pi}=(\vec{\nu}_{0},\vec{\nu}_{1},\ldots)\in\vec{\Pi}$ by
$\vec{\pi}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\ldots,\mathbf{x}_{k};u_{0},\ldots,u_{k-1}):=\mu_{k+1}({\mathrm{d}}u_{k}|x,w,x^{2}_{1},\ldots,x^{k+1}_{k};u,u_{0},\ldots,u_{k-1})$
for each $k\in{\mathbb{N}}$. Then, similarly as in the proof of Claim 2, we have
$J_{\pi^{1}}(1;x,w;u)$ $=\sum_{k=1}^{\infty}\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{k}(x,w,x_{2},\ldots,x_{k};u,u_{1},\ldots,u_{k-1},u_{k})$
$\qquad\mu_{k}({\mathrm{d}}u_{k}|x,w,x_{2},\ldots,x_{k};u,u_{1},\ldots,u_{k-1})$
$\qquad p_{k-1}({\mathrm{d}}x_{k}|x,w,x_{2},\ldots,x_{k-1};u,u_{1},\ldots,u_{k-1})\cdots$
$\qquad\mu_{2}({\mathrm{d}}u_{2}|x,w,x_{2};u,u_{1})\,p_{1}({\mathrm{d}}x_{2}|x,w;u,u_{1})\,\mu_{1}({\mathrm{d}}u_{1}|x,w;u)$
$=\sum_{k=1}^{\infty}\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{k}(x,w,x^{2}_{1},\ldots,x^{k}_{k-1};u,u_{0},\ldots,u_{k-2},u_{k-1})$
$\qquad\mu_{k}({\mathrm{d}}u_{k-1}|x,w,x^{2}_{1},\ldots,x^{k}_{k-1};u,u_{0},\ldots,u_{k-2})$
$\qquad p_{k-1}({\mathrm{d}}x^{k}_{k-1}|x,w,x^{2}_{1},\ldots,x^{k-1}_{k-2};u,u_{0},\ldots,u_{k-2})\cdots$
$\qquad\mu_{2}({\mathrm{d}}u_{1}|x,w,x^{2}_{1};u,u_{0})\,p_{1}({\mathrm{d}}x^{2}_{1}|x,w;u,u_{0})\,\mu_{1}({\mathrm{d}}u_{0}|x,w;u)$
$=\sum_{k=0}^{\infty}\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{k+1}(x,w,x^{2}_{1},\ldots,x^{k+1}_{k};u,u_{0},\ldots,u_{k-1},u_{k})$
$\qquad\mu_{k+1}({\mathrm{d}}u_{k}|x,w,x^{2}_{1},\ldots,x^{k+1}_{k};u,u_{0},\ldots,u_{k-1})$
$\qquad p_{k}({\mathrm{d}}x^{k+1}_{k}|x,w,x^{2}_{1},\ldots,x^{k}_{k-1};u,u_{0},\ldots,u_{k-1})\cdots$
$\qquad\mu_{2}({\mathrm{d}}u_{1}|x,w,x^{2}_{1};u,u_{0})\,p_{1}({\mathrm{d}}x^{2}_{1}|x,w;u,u_{0})\,\mu_{1}({\mathrm{d}}u_{0}|x,w;u)$
$=\mathbf{J}_{\vec{\pi}}(1;x,w,x,x,\ldots;u,u_{\Delta},u_{\Delta},\ldots).$
This yields Claim 4.
Thanks to the four established claims above, we have (A.13). This concludes the proof of Theorem A.3. ∎
A.4. Proof of Theorem A.4.
We will make use of facts established in section A.3. Let $u_{\Delta}\in U$. Recall that, by (A.11), we have
(A.15) $\mathbf{J}^{\ast}(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots)=\inf_{\vec{\pi}\in\vec{\Pi}^{\prime}}\mathbf{J}_{\vec{\pi}}(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots).$
Since $\mathbf{g}$ is non-negative, the infimum in the right-hand side of (A.15) can, according to Proposition 9.19 of [BS], be taken over non-randomized policies $\vec{\pi}\in\vec{\Pi}^{\prime}$, i.e., over every $\vec{\pi}=(\vec{\mu}_{0},\vec{\mu}_{1},\ldots)\in\vec{\Pi}^{\prime}$ such that, for each $k\in{\mathbb{N}}_{0}$, $\mathbf{x}_{0}$, $\ldots$, $\mathbf{x}_{k}\in\mathbf{S}$, and $u_{0}$, $\ldots$, $u_{k-1}\in U$, the measure $\vec{\mu}_{k}({\mathrm{d}}u_{k}|\mathbf{x}_{0},\,\ldots,\,\mathbf{x}_{k};\,u_{0},\,\ldots,u_{k-1})$ is a Dirac measure.
Next, following the proof of Claim 1 in section A.3, we can deduce via defining a corresponding non-randomized policy in $\Pi$ (see Theorem A.4 for its definition) for any non-randomized policy in $\vec{\Pi}^{\prime}$ that
(A.16) $\mathbf{J}^{\ast}(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots)\geq\inf_{\pi\in\Pi}J_{\pi}(x).$
Similarly, following the proof of Claim 3 in section A.3, we can see that
(A.17) $\mathbf{J}^{\ast}(0;x,x,\ldots;u_{\Delta},u_{\Delta},\ldots)\leq\inf_{\pi\in\Pi}J_{\pi}(x).$
Hence, by the first line of (A.13) together with (A.16) and (A.17), we immediately have (A.8). This concludes the proof of Theorem A.4. ∎
Appendix B Some technical proofs
B.1. Proof of Lemma 4.4
First, we prove the auxiliary Lemmas B.1, B.2, B.3.
Lemma B.1.
Consider the maps $\iota_{i}(s,x)\mapsto\iota_{i}(s,x)$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, $i\in\{1,2\}$, and $\iota_{3}:(s,x,\alpha)\mapsto\iota_{3}(s,x,\alpha)$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}_{+}\times\Omega$, respectively defined by
$[\iota_{1}(s,x,\alpha)](t)$ $:=\int_{0}^{t}f(s+r,x(\cdot\wedge s),\alpha(s+r))\,{\mathrm{d}}r,\quad t\in{\mathbb{R}}_{+},$
$[\iota_{2}(s,x,\alpha)](t)$ $:=\iota_{2}^{t}(s,x,\alpha),\quad t\in{\mathbb{R}}_{+},$
$[\iota_{3}(s,x,\alpha)](t)$ $:=(s,x(t\wedge s),\alpha),\quad t\in{\mathbb{R}}_{+},$
where $\iota_{2}^{t}$ is defined by (B.1) below. We have that
  • (i)
    the map $\iota_{1}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$;
  • (ii)
    the map $\iota_{2}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$;
  • (iii)
    the map $\iota_{3}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$.
Proof of Lemma B.1. Item (i) directly follows from (ii)-(iii), noticing that $\iota_{1}=\iota_{2}\circ\iota_{3}$. Item (iii) is an adaption of Theorem 96 (b), 146-IV, in [DMprobPot]. In order to prove item (ii), we consider the map
(B.1) $\iota_{2}^{t}:(s,x,\alpha)\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r,\quad{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}^{d}.$
The proof follows the line of the reasoning on pp. 84 and 135 in [Billingsley] regarding the measurability of path-space valued random elements, using that
(ii’) for any $t\in{\mathbb{R}}_{+}$, the map $\iota_{2}^{t}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to $\mathcal{B}({\mathbb{R}}^{d})$,
which we show next. First note that the maps $s\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r=\int_{s}^{s+t}f(r,x,\alpha(r))\,{\mathrm{d}}r$, ${\mathbb{R}}_{+}\to{\mathbb{R}}^{d}$, $(x,\alpha)\in\Omega\times\mathcal{A}$, are continuous. Moreover, for every $s\in{\mathbb{R}}_{+}$, the map $(x,\alpha)\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r$, $\Omega\times\mathcal{A}\to{\mathbb{R}}^{d}$, is ${\mathcal{F}}^{0}_{t+s}\otimes\mathcal{B}(\mathcal{A})$- and hence ${\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$-measurable.171717 To see this, consider the case that each component $f^{j}$ of $f=(f^{1},\ldots,f^{d})$ is of the form $f^{j}(r,x,a)=f^{j}_{1}(x(\cdot\wedge(t+s)))\,f^{j}_{2}(r,a)$, where $f^{j}_{1}$ is ${\mathcal{F}}^{0}$- and $f^{j}_{2}$ is (+)×$\mathcal{B}({\mathbb{R}}_{+})\times\mathcal{B}(A)$-measurable. Taking Remark 3.1 into account, we can deduce the ${\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$-measurability of $(x,\alpha)\mapsto\int_{0}^{t}f(s+r,x,\alpha(s+r))\,{\mathrm{d}}r$. It remains to apply a monotone-class argument. The reasoning here is very similar to the proof of Lemma 2 in [Yushkevich80]. These two facts establish (ii’) (see  Proposition 1.13, p. 5, in [KaratzasShreve]). ∎
Lemma B.2.
The map $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(0)}$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to\Omega$, defined by
$\phi^{s,x,\alpha}_{(0)}(t):=x(t).{\mathbf{1}}_{[0,s)}(t)+\Bigl{[}x(t)+\int_{s}^{t}f(r,x(\cdot\wedge s),\alpha(r))\,{\mathrm{d}}r\Bigr{]}.{\mathbf{1}}_{[s,\infty)}(t),\quad t\in{\mathbb{R}}_{+},$
is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$.
Proof of Lemma B.2. By Theorem 96 (d), 146-IV, in [DMprobPot], the map $\iota_{4}:(s,x,\tilde{x})\mapsto\iota_{4}(s,x)$, +×${\mathbb{R}}_{+}\times\Omega\times\Omega\to\Omega$, defined by
$[\iota_{4}(s,x,\tilde{x})](t):=x(t).{\mathbf{1}}_{[0,s)}(t)+\tilde{x}(t-s).{\mathbf{1}}_{[s,\infty)}(t),\quad t\in{\mathbb{R}}_{+},$
is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes{\mathcal{F}}^{0}$ to ${\mathcal{F}}^{0}$. This allows to conclude the proof, being $\phi^{s,x,\alpha}_{(0)}=\iota_{4}(s,x,\iota_{1}(s,x,\alpha))$, and recalling by Lemma B.1-(i) the measurability of $(s,x,\alpha)\mapsto(s,x,\iota_{1}(s,x,\alpha))$, +×${\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}_{+}\times\Omega$, from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes{\mathcal{F}}^{0}$. ∎
Lemma B.3.
Consider the mappings $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(n)}$, +×${\mathbb{R}}_{+}\times\Omega\to\Omega$, $n\in{\mathbb{N}}$, recursively defined by
$\phi^{s,x,\alpha}_{(n+1)}(t):=x(t).{\mathbf{1}}_{[0,s)}(t)+\Bigl{[}x(t)+\int_{s}^{t}f(r,\phi^{s,x}_{\text{($n$)}},\alpha(r))\,{\mathrm{d}}r\Bigr{]}.{\mathbf{1}}_{[s,\infty)}(t),\quad t\in{\mathbb{R}}_{+}.$
Then, for every $n\in{\mathbb{N}}_{0}$, the map $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(n)}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$.
Proof of Lemma B.3. Fix $n\in{\mathbb{N}}_{0}$. Assume that $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(n)}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$. To show that $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(n+1)}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$, one can proceed essentially as in the proof of Lemma B.2 but one should use Lemma B.1-(ii) instead of Lemma B.1-(i) and note that
$\phi^{s,x,\alpha}_{(n+1)}=\iota_{4}(s,x,\iota_{2}(s,\phi^{s,x,\alpha}_{(n)},\alpha)).$
Finally, note that, by Lemma B.2, $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(0)}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$. The proof is concluded by mathematical induction. ∎
We can finally prove Lemma 4.4.
Proof of Lemma 4.4. By Lemma B.3, for every $t\in{\mathbb{R}}_{+}$, the function
$(s,x,\alpha)\mapsto\phi^{s,x,\alpha}(t),\,\,{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}^{d},$
is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to $\mathcal{B}({\mathbb{R}}^{d})$ as the limit of the functions
$(s,x,\alpha)\mapsto\phi^{s,x,\alpha}_{(n)}(t),\,\,{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}\to{\mathbb{R}}^{d},n\in{\mathbb{N}},$
which are all measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to $\mathcal{B}({\mathbb{R}}^{d})$.
Moreover, for each $(s,x,\alpha)\in{\mathbb{R}}_{+}\times\Omega\times\mathcal{A}$, the function $t\mapsto\phi^{s,x,\alpha}(t)$, ${\mathbb{R}}_{+}\to{\mathbb{R}}^{d}$, is right-continuous. Thus, following the proof of Proposition 1.13 on p. 5 in [KaratzasShreve], one can show that
$(s,x,\alpha,t)\mapsto\phi^{s,x,\alpha}(t),\,\,{\mathbb{R}}_{+}\times\Omega\times A\times{\mathbb{R}}_{+}\to{\mathbb{R}}_{+},$
is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})\otimes\mathcal{B}({\mathbb{R}}_{+})$ to $\mathcal{B}({\mathbb{R}}^{d})$. Hence, for every $t\in{\mathbb{R}}_{+}$,
$(s,x,\alpha)\mapsto\phi^{s,x,\alpha}(t),\,\,{\mathbb{R}}_{+}\times\Omega\to{\mathbb{R}}^{d},$
is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to $\mathcal{B}({\mathbb{R}}^{d})$, from which we finally can conclude, exactly as in the proof of Lemma B.1-(ii), that $(s,x,\alpha)\mapsto\phi^{s,x,\alpha}$ is measurable from $\mathcal{B}({\mathbb{R}}_{+})\otimes{\mathcal{F}}^{0}\otimes\mathcal{B}(\mathcal{A})$ to ${\mathcal{F}}^{0}$. ∎
B.2. Proof of Lemma 4.15
Let
$y:=\phi^{s,x,\alpha}\otimes_{t}e$
and note that, for every $\mathbf{a}\in\mathbb{A}^{\prime}$, we can deduce in the same way as (4.8) that
(B.2) $J(t,y,\mathbf{a})=\sum_{n=0}^{\infty}\check{{{\mathbb{E}}}}^{t,y,\mathbf{a}}_{n+1}\left[\vec{\ell^{1}_{n}}\right],$
where the functions $\vec{\ell_{n}^{1}}:\check{\Omega}_{n+1}\to{\mathbb{R}}_{+}$, $n\in{\mathbb{N}}_{0}$, are defined by
(B.3) $\begin{split}\vec{\ell_{n}^{1}}(\check{\omega}_{n+1})&:=\int_{t_{n}\wedge T}^{T}\chi^{t_{n},X^{t,y}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(\tau)\,\ell(t,X^{t,y}_{n}(\check{\omega}_{n+1}),\alpha_{n}(\tau))\,{\mathrm{d}}\tau\\ &\qquad+{\mathbf{1}}_{[0,T]}(t_{n})\,\chi^{t_{n},X^{t,y}_{n}(\check{\omega}_{n+1}),\alpha_{n}}(T)\,h(X^{t,y}_{n}(\check{\omega}_{n+1})).\end{split}$
for each $\check{\omega}_{n+1}=(t_{0},e_{0},\alpha_{0};\ldots;t_{n},e_{n},\alpha_{n})\in\check{\Omega}_{n+1}$.
In the rest of this proof, we use the marginals $r^{1}_{N}(\pi^{1},p^{1})$ on $(S\times U)^{N}$, $N\in\{2,3,\ldots\}$ defined by (A.4). We want to emphasize that we still use the data specified earlier in section 4.5. We also use the space $\mathbb{A}^{1}$ of all policies $\mathbf{a}^{1}=(\mathbf{a}^{1}_{1},\mathbf{a}^{1}_{2},\ldots)$ such that each $\mathbf{a}^{1}_{j}=\mathbf{a}^{1}_{j}(d\alpha_{j}|(t_{0},e_{0}),\ldots,(t_{j},e_{j});\alpha_{0},\ldots,\alpha_{j-1})$, $j\in{\mathbb{N}}$, is a universally measurable stochastic kernel on $U$ given $S^{j+1}\times U^{j}$.
First note that, for every $\mathbf{a}^{1}\in\mathbb{A}^{1}$,
$\int g_{1}\,{\mathrm{d}}r^{1}_{2}(\mathbf{a}^{1},\delta_{((s,x(s)),(t,e);\alpha)})$ $=\int_{S^{2}\times U}\int_{U}g_{1}((t_{0},e_{0}),(t_{1},e_{1});\alpha_{0},\alpha_{1})$
$\qquad\qquad\mathbf{a}^{1}_{1}({\mathrm{d}}\alpha_{1}|(t_{0},e_{0}),(t_{1},e_{1});\alpha_{0})$
$\qquad\qquad\delta_{((s,x(s)),(t,e);\alpha)}({\mathrm{d}}t_{0}\,{\mathrm{d}}e_{0}\,{\mathrm{d}}t_{1}\,{\mathrm{d}}e_{1}\,{\mathrm{d}}\alpha_{0})$
$=\int_{U}\vec{\ell}_{1}((s,x(s),\alpha),(t,e,\alpha_{1}))$
$\qquad\qquad\mathbf{a}^{1}_{1}({\mathrm{d}}\alpha_{1}|(s,x(s)),(t,e);\alpha).$
By the definition of $\vec{\ell}_{1}$ in (4.9), the terms $\vec{\ell}_{1}((s,x(s),\alpha),(t,e,\alpha_{1}))$ depend on
[…]
Together with the definition of $\vec{\ell_{0}^{1}}$ in (B.3), we obtain
$\int g_{1}\,{\mathrm{d}}r^{1}_{2}(\mathbf{a}^{1},\delta_{((s,x(s)),(t,e);\alpha)})$
$=\int_{U}\vec{\ell_{0}^{1}}(t,y(t),\alpha_{1})\,\mathbf{a}^{1}_{1}({\mathrm{d}}\alpha_{1}|(s,x(s)),(t,e);\alpha)$
$=\int_{U}\vec{\ell_{0}^{1}}(t,y(t),\alpha_{0})\,\mathbf{a}^{1}_{1}({\mathrm{d}}\alpha_{0}|(s,x(s)),(t,e);\alpha)$
=(4.4)$\check{{{\mathbb{E}}}}_{1}^{t,y,\mathbf{a}}\left[\vec{\ell_{0}^{1}}\right]$
with $\mathbf{a}=(\mathbf{a}_{0},\mathbf{a}_{1},\ldots)\in\mathbb{A}^{\prime}$ such that
$\mathbf{a}_{0}({\mathrm{d}}\alpha_{0}|t_{0},e_{0})=\mathbf{a}^{1}_{1}({\mathrm{d}}\alpha_{0}|(s,x(s)),(t_{0},e_{0});\alpha).$
Next, let $N\in\{2,3,\ldots\}$ and $\mathbf{a}^{1}\in\mathbb{A}^{1}$. Then
$\int g_{N}\,{\mathrm{d}}r^{1}_{N+1}(\mathbf{a}^{1},\delta_{((s,x(s)),(t,e);\alpha)})$
$=\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}g_{N}((s,x(s)),(t,e),(t_{2},e_{2}),\ldots,(t_{N},e_{N});\alpha,\alpha_{1},\ldots,\alpha_{N})$
$\qquad\mathbf{a}^{1}_{N}({\mathrm{d}}\alpha_{N}|(s,x(s)),(t,e),(t_{2},e_{2}),\ldots,(t_{N},e_{N});\alpha,\alpha_{1},\ldots,\alpha_{N-1})$
$\qquad\vec{Q}_{N-1}({\mathrm{d}}t_{N}\,{\mathrm{d}}e_{N}|(s,x(s)),(t,e),(t_{2},e_{2})\,\ldots,(t_{N-1},e_{N-1});\alpha,\alpha_{1},\ldots,\alpha_{N-1})\cdots$
$\qquad\mathbf{a}^{1}_{2}({\mathrm{d}}\alpha_{2}|(s,x(s)),(t,e),(t_{2},e_{2});\alpha,\alpha_{1})$
$\qquad\vec{Q}_{1}({\mathrm{d}}t_{2}\,{\mathrm{d}}e_{2}|(s,x(s)),(t,e);\alpha,\alpha_{1})\,\mathbf{a}^{1}_{1}({\mathrm{d}}\alpha_{1}|(s,x(s)),(t,e);\alpha).$
Since (up to an $r^{1}_{N+1}(\mathbf{a}^{1},\delta_{((s,x(s)),(t,e);\alpha)})$-null set)
(B.4) $\begin{split}&g_{N}((s,x(s)),(t,e),(t_{2},e_{2}),\ldots,(t_{N},e_{N});\alpha,\alpha_{1},\ldots,\alpha_{N})\\ &\qquad=\vec{\ell}_{N}((s,x(s),\alpha),(t,e,\alpha_{1}),(t_{2},e_{2},\alpha_{2}),\ldots,(t_{N},e_{N},\alpha_{N}))\end{split}$
and the right-hand side of (B.4) depends on
$X^{s,x}_{N}(s,x(s),\alpha;t,e,\alpha_{1};t_{2},e_{2},\alpha_{2};\ldots;t_{N},e_{N},\alpha_{N}),$
which equals
$X^{t,y}_{N-1}(t,y(t),\alpha_{1};t_{2},e_{2},\alpha_{2};\ldots;t_{N},e_{N},\alpha_{N}),$
we have (again up to an $r^{1}_{N+1}(\mathbf{a}^{1},\delta_{((s,x(s)),(t,e);\alpha)})$-null set)
$g_{N}((s,x(s)),(t,e),(t_{2},e_{2}),\ldots,(t_{N},e_{N});\alpha,\alpha_{1},\ldots,\alpha_{N})$
$\qquad=\overrightarrow{\ell^{1}_{N-1}}((t,y(t),\alpha_{1});(t_{2},e_{2},\alpha_{2});\ldots;(t_{N},e_{N},\alpha_{N})).$
Thus
$\int g_{N}\,{\mathrm{d}}r^{1}_{N+1}(\mathbf{a}^{1},\delta_{((s,x(s)),(t,e);\alpha)})$
$=\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}\overrightarrow{\ell^{1}_{N-1}}((t,y(t),\alpha_{1});(t_{2},e_{2},\alpha_{2});\ldots;(t_{N},e_{N},\alpha_{N}))$
$\,\,\,\,\mathbf{a}^{1}_{N}({\mathrm{d}}\alpha_{N}|(s,x(s)),(t,y(t)),(t_{2},e_{2}),\ldots,(t_{N},e_{N});\alpha,\alpha_{1},\ldots,\alpha_{N-1})$
$\,\,\,\,Q(t_{N},{\mathrm{d}}e_{N}|t_{N-1},\underbrace{X^{s,x}_{N-1}(s,x(s),\alpha;t,e,\alpha_{1};t_{2},e_{2},\alpha_{2};\ldots;t_{N-1},e_{N-1},\alpha_{N-1})}_{=X^{t,y}_{N-2}(t,y(t),\alpha_{1};t_{2},e_{2},\alpha_{2};\ldots;t_{N-1},e_{N-1},\alpha_{N-1})},\alpha_{N-1})$
$\,\,\,\,[-{\mathrm{d}}F(t_{N}|t_{N-1},X^{t,y}_{N-2}(t,y(t),\alpha_{1};t_{2},e_{2},\alpha_{2};\ldots;t_{N-1},e_{N-1},\alpha_{N-1}),\alpha_{N-1})]\cdots$
$\,\,\,\,\mathbf{a}^{1}_{2}({\mathrm{d}}\alpha_{2}|(s,x(s)),(t,y(t)),(t_{2},e_{2});\alpha,\alpha_{1})$
$\,\,\,\,\underbrace{Q(t_{2},{\mathrm{d}}e_{2}|t,\overbrace{X^{s,x}_{1}(s,x(s),\alpha;t,e,\alpha_{1})}^{=X^{t,y}_{0}(t,y(t),\alpha_{1})=\phi^{t,y,\alpha_{1}}},\alpha_{1})}_{=Q^{t,y,\alpha_{1}}(t_{2},{\mathrm{d}}e_{2})}\,[-{\mathrm{d}}F^{t,y,\alpha_{1}}(t_{2})]$
$\,\,\,\,\mathbf{a}_{1}^{1}({\mathrm{d}}\alpha_{1}|(s,x(s)),(t,y(t));\alpha)$
$=^{(\text{after relabeling of variables})}$
$\,\,\,\,\int_{U}\int_{S}\int_{U}\cdots\int_{S}\int_{U}\overrightarrow{\ell^{1}_{N-1}}((t,y(t),\alpha_{0});(t_{1},e_{1},\alpha_{1});\ldots;(t_{N-1},e_{N-1},\alpha_{N-1}))$
$\,\,\,\,\mathbf{a}^{1}_{N}({\mathrm{d}}\alpha_{N-1}|(s,x(s)),(t,y(t)),(t_{1},e_{1}),\ldots,(t_{N-1},e_{N-1});\alpha,\alpha_{0},\ldots,\alpha_{N-2})$
$\,\,\,\,Q(t_{N-1},{\mathrm{d}}e_{N-1}|t_{N-2},X^{t,y}_{N-2}(t,y(t),\alpha_{0};t_{1},e_{1},\alpha_{1};\ldots;t_{N-2},e_{N-2},\alpha_{N-2}),\alpha_{N-2})$
$\,\,\,\,[-{\mathrm{d}}F(t_{N-1}|t_{N-2},X^{t,y}_{N-2}(t,y(t),\alpha_{0};t_{1},e_{1},\alpha_{1};\ldots;t_{N-2},e_{N-2},\alpha_{N-2}),\alpha_{N-2})]\cdots$
$\,\,\,\,\mathbf{a}^{1}_{2}({\mathrm{d}}\alpha_{1}|(s,x(s)),(t,y(t)),(t_{1},e_{1});\alpha,\alpha_{0})\,Q^{t,y,\alpha_{0}}(t_{1},{\mathrm{d}}e_{1})\,[-{\mathrm{d}}F^{t,y,\alpha_{0}}(t_{1})]$
$\,\,\,\,\mathbf{a}_{1}^{1}({\mathrm{d}}\alpha_{0}|(s,x(s)),(t,y(t));\alpha)$
$=\check{{{\mathbb{E}}}}_{N}^{t,y,\mathbf{a}}\Bigl{[}\overrightarrow{\ell^{1}_{N-1}}\Bigr{]}$
with $\mathbf{a}=(\mathbf{a}_{0},\mathbf{a}_{1},\ldots)\in\mathbb{A}^{\prime}$ such that, for each $n\in{\mathbb{N}}$,
$\mathbf{a}_{n-1}({\mathrm{d}}\alpha_{N-1}|(t_{0},e_{0}),\ldots,(t_{N-1},e_{N-1});\alpha_{0},\ldots,\alpha_{N-1})$
$\qquad=\mathbf{a}^{1}_{n}({\mathrm{d}}\alpha_{N-1}|(s,x(s)),(t_{0},e_{0}),\ldots,(t_{N-1},e_{N-1});\alpha,\alpha_{0},\ldots,\alpha_{N-1}).$
Similarly, one can show that, for every $\mathbf{a}\in\mathbb{A}^{\prime}$, there exists an $\mathbf{a}^{1}\in\mathbb{A}^{1}$ such that
$\check{{{\mathbb{E}}}}_{N}^{t,y,\mathbf{a}}\Bigl{[}\overrightarrow{\ell^{1}_{N-1}}\Bigr{]}=\int g_{N}\,{\mathrm{d}}r^{1}_{N+1}(\mathbf{a}^{1},\delta_{((s,x(s)),(t,e);\alpha)}).$
Consequently, we have
[…]
$=\inf_{\mathbf{a}^{1}\in\mathbb{A}^{1}}\sum_{n=0}^{\infty}\int g_{n+1}\,{\mathrm{d}}r^{1}_{n+2}(\mathbf{a}^{1},\delta_{((s,x(s)),(t,e);\alpha)})$
=(A.5)$\inf_{\mathbf{a}^{1}\in\mathbb{A}^{1}}J_{\mathbf{a}^{1}}(1;(s,x(s)),(t,e);\alpha)$ )
=(A.6)$J^{\ast}(1;(s,x(s)),(t,e);\alpha).$ ) .
This concludes Case 2 and thus the proof. ∎
B.3. Proof of Lemma 5.6
Without loss of generality, let $s_{1}=0$ and $s_{2}=T$.
Fix $s\in[0,T]$, $x$, $\tilde{x}\in\Omega$, $a\in\mathcal{A}$, and $e\in{\mathbb{R}}^{d}$. Put
$(\phi,\ell,\lambda,\chi,Q)$ $:=(\phi^{s,x,a},\ell^{s,x,a},\lambda^{s,x,a},\chi^{s,x,a},Q^{s,x,a}),$
$(\tilde{\phi},\tilde{\ell},\tilde{\lambda},\tilde{\chi},\tilde{Q})$ $:=(\phi^{s,\tilde{x},a},\ell^{s,\tilde{x},a},\lambda^{s,\tilde{x},a},\chi^{s,\tilde{x},a},Q^{s,\tilde{x},a}).$
By Assumption 4.1, for all $t\in[s,T]$,
[…]
Let us set
$[(G_{0,T;\eta,a})\psi](s,x)-[(G_{0,T;\eta,a})\psi](s,\tilde{x})=I_{1}+I_{2}+I_{3},$
where
$I_{1}$ $:=\chi(T)\,[\eta(T,\phi)-\eta(T,\tilde{\phi})]+[\chi(T)-\tilde{\chi}(T)]\,\eta(T,\tilde{\phi}),$
$I_{2}$ $:=\int_{s}^{T}\chi(t)[\ell(t)-\tilde{\ell}(t)]+[\chi(t)-\tilde{\chi}(t)]\tilde{\ell}(t)\,{\mathrm{d}}t,$
$I_{3}$ $:=J_{1}+J_{2}+J_{3}+J_{4},$
with
$J_{1}$ $:=\int_{s}^{T}\Bigl{[}\int_{{\mathbb{R}}^{d}}\psi(t,\phi\otimes_{t}e)\lambda(t)\chi(t)\,Q(t,{\mathrm{d}}e)-\int_{{\mathbb{R}}^{d}}\psi(t,\phi\otimes_{t}e)\lambda(t)\chi(t)\,\tilde{Q}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t,$
$J_{2}$ $:=\int_{s}^{T}\Bigl{[}\int_{{\mathbb{R}}^{d}}\psi(t,\phi\otimes_{t}e)\lambda(t)[\chi(t)-\tilde{\chi}(t)]\,\tilde{Q}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t,$
$J_{3}$ $:=\int_{s}^{T}\Bigl{[}\int_{{\mathbb{R}}^{d}}\psi(t,\phi\otimes_{t}e)[\lambda(t)-\tilde{\lambda}(t)]\tilde{\chi}(t)\,\tilde{Q}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t,$
$J_{4}$ $:=\int_{s}^{T}\Bigl{[}\int_{{\mathbb{R}}^{d}}[\psi(t,\phi\otimes_{t}e)-\psi(t,\tilde{\phi}\otimes_{t}e)]\tilde{\lambda}(t)\tilde{\chi}(t)\,\tilde{Q}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t.$
Then,
$I_{1}$ $=\chi(T)\,[\eta(T,\phi)-\eta(T,\tilde{\phi})]+[\chi(T)-\tilde{\chi}(T)]\,\eta(T,\tilde{\phi})$
$\leq c^{\prime}{\mathrm{e}}^{L_{f}(T-s)}\lVert{x-\tilde{x}}\rVert_{s}+L_{f}(T-s){\mathrm{e}}^{L_{f}(T-s)}\lVert{x-\tilde{x}}\rVert_{s}\,\lVert{\eta}\rVert_{\infty}$
$\leq\left[c^{\prime}{\mathrm{e}}^{L_{f}T}+(T-s)\check{L}\right]\lVert{x-\tilde{x}}\rVert_{s}$
thanks to (B.5) and Assumption 4.1, and
$I_{2}$ $=\int_{s}^{T}\chi(t)[\ell(t)-\tilde{\ell}(t)]+[\chi(t)-\tilde{\chi}(t)]\tilde{\ell}(t)\,{\mathrm{d}}t$
$\leq L_{f}(T-s){\mathrm{e}}^{L_{f}(T-s)}\lVert{x-\tilde{x}}\rVert_{s}+C_{f}(T-s){\mathrm{e}}^{L_{f}(T-s)}\lVert{x-\tilde{x}}\rVert_{s}$
$\leq(T-s)\check{L}\lVert{x-\tilde{x}}\rVert_{s}$
thanks to (B.5) and Assumption 4.1.
Moreover,
$J_{1}$ $=\int_{s}^{T}\Bigl{[}\int_{{\mathbb{R}}^{d}}\psi(t,\phi\otimes_{t}e)\lambda(t)\chi(t)\,Q(t,{\mathrm{d}}e)-\int_{{\mathbb{R}}^{d}}\psi(t,\phi\otimes_{t}e)\lambda(t)\chi(t)\,\tilde{Q}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t$
$\leq C_{\lambda}(T-s)L_{Q}\,c\,{\mathrm{e}}^{L_{f}(T-s)}\,\lVert{x-\tilde{x}}\rVert_{s}$
$\leq(T-s)\check{L}\,c\,\lVert{x-\tilde{x}}\rVert_{s}$
thanks to Assumptions 4.1 and 4.2 together with (5.5) and (B.5),
$J_{2}$ $=\int_{s}^{T}\Bigl{[}\int_{{\mathbb{R}}^{d}}\psi(t,\phi\otimes_{t}e)\lambda(t)[\chi(t)-\tilde{\chi}(t)]\,\tilde{Q}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t$
$\leq(T-s)\lVert{\psi}\rVert_{\infty}\,C_{\lambda}\,L_{f}(T-s){\mathrm{e}}^{L_{f}(T-s)}\lVert{x-\tilde{x}}\rVert_{s}$
$\leq(T-s)\check{L}\,\lVert{\psi}\rVert_{\infty}\,\lVert{x-\tilde{x}}\rVert_{s}$
thanks to (B.5) and Assumption 4.1,
$J_{3}$ $=\int_{s}^{T}\Bigl{[}\int_{{\mathbb{R}}^{d}}\psi(t,\phi\otimes_{t}e)[\lambda(t)-\tilde{\lambda}(t)]\tilde{\chi}(t)\,\tilde{Q}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t$
$\leq(T-s)\lVert{\psi}\rVert_{\infty}\,L_{f}{\mathrm{e}}^{L_{f}(T-s)}\,\lVert{x-\tilde{x}}\rVert_{s}$
$\leq(T-s)\check{L}\,\lVert{\psi}\rVert_{\infty}\,\lVert{x-\tilde{x}}\rVert_{s}$
thanks to (B.5) and Assumption 4.1, and
$J_{4}$ $=\int_{s}^{T}\Bigl{[}\int_{{\mathbb{R}}^{d}}[\psi(t,\phi\otimes_{t}e)-\psi(t,\tilde{\phi}\otimes_{t}e)]\tilde{\lambda}(t)\tilde{\chi}(t)\,\tilde{Q}(t,{\mathrm{d}}e)\Bigr{]}\,{\mathrm{d}}t$
$\leq(T-s)\,c\,{\mathrm{e}}^{L_{f}(T-s)}\,\lVert{x-\tilde{x}}\rVert_{s}\,C_{\lambda}$
$\leq(T-s)\check{L}\,c\,\lVert{x-\tilde{x}}\rVert_{s}$
thanks to (5.5), (B.5) and Assumption 4.1. Therefore we can see that (LABEL:E:GpsiStab:xe) holds. ∎
B.4. Proof of Lemma 5.7
First, given $c^{\prime}\in{\mathbb{R}}_{+}$, define $(c_{n}(c^{\prime}))_{n\in{\mathbb{N}}_{0}}$ recursively by $c_{0}(c^{\prime})=0$ and
$c_{n+1}(c^{\prime}):=\left(c^{\prime}{\mathrm{e}}^{L_{f}T}+\frac{1}{2}\right)+\frac{1}{2}c_{n}(c^{\prime}).$
By mathematical induction,
(B.6) $c_{n}(c^{\prime})\leq 2c^{\prime}{\mathrm{e}}^{L_{f}T}+1\text{ for all $n\in{\mathbb{N}}$.}$
Next, recall the partition $0=r_{0}<r_{1}<\ldots<r_{N}=T$ from Step 1 of the proof of Theorem 5.4 (i). Note that $r_{j+1}-r_{j}<\frac{1}{2}$ for all $j$.
Let $\psi_{0}\equiv 0$ and consider a subpartition $(s_{j,k})_{j=0,\ldots,{N-1};k=0,\ldots,M}$ with a sufficiently large integer $M$ such that, for each $j\in\{1,\ldots,N\}$, we have
$r_{j-1}=s_{j-1,0}<s_{j-1,1}<\ldots<s_{j-1,M}=r_{j}$
and, for each $k\in\{0,\ldots,M-1\}$,
(B.7) $s_{j,k+1}-s_{j,k}\leq\frac{1}{12\check{L}{\color[rgb]{0,0,0}\left(1+\sup_{n\in{\mathbb{N}}_{0}}\lVert{(G^{n}_{r_{j-1},r_{j};\bar{\psi}})\psi_{0}}\rVert_{\infty}\right)}}=:\check{m}_{j}.$
Step 1. Let $n\in{\mathbb{N}}_{0}$. By (B.7) and Lemma 5.6 applied to $[(G^{n}_{r_{N-1},r_{N};\bar{\psi}})\psi_{0}]$ restricted to $[s_{N-1,M-1},s_{N-1,M}]\times\Omega$, $c_{n}(L_{f})$, $L_{f}$, $h$, $s_{N-1,M-1}$, $s_{N-1,M}$ in place of $\psi$, $c$, $c^{\prime}$, $\eta$, $s_{1}$, $s_{2}$ and noting that
$(G_{s_{N-1,M-1},s_{N-1,M};h})(\psi_{0}|_{[s_{N-1,M-1},s_{N-1,M}]\times\Omega})$
$\qquad=[(G_{r_{N-1},r_{N};\bar{\psi}})\psi_{0})|_{[s_{N-1,M-1},s_{N-1,M}]\times\Omega},$
we can see that
$\left|{[G_{r_{N-1},r_{N};\bar{\psi}}^{n+1}\psi_{0}](s,x)-[G_{r_{N-1},r_{N};\bar{\psi}}^{n+1}\psi_{0}](s,\tilde{x})}\right|\leq\left(L_{f}{\mathrm{e}}^{L_{f}T}+\frac{1+c_{n}(L_{f})}{2}\right)\lVert{x-\tilde{x}}\rVert_{s}$
[…]
for all $s\in[s_{N-1,M-1},s_{N-1,M}]$, $x$, $\tilde{x}\in\Omega$.
By Theorem 5.4 (ii), letting $n\to\infty$ yields (5.7) on $[s_{N-1,M-1},s_{N-1,M}]\times\Omega\times\Omega$ with $c^{\prime}_{N}:=2L_{f}{\mathrm{e}}^{L_{f}T}+1$ in place of $L_{\bar{\psi}}$.
Step 2. Let $n\in{\mathbb{N}}_{0}$. Without loss of generality, assume that
$r_{N-1}\leq s_{N-1,M-2}.$
Consequently, by (B.7) and Lemma 5.6 applied to $[(G^{n}_{r_{N-1},r_{N};\bar{\psi}})\psi]$ restricted to $[s_{N-1,M-2},s_{N-1,M-1}]\times\Omega$, $c_{n}(c^{\prime}_{N})$, $c^{\prime}_{N}$, $\bar{\psi}$, $s_{N-1,M-2}$, $s_{N-1,M-1}$ in place of $\psi$, $c$, $c^{\prime}$, $\eta$, $s_{1}$, $s_{2}$, we can with Step 1 of this proof deduce that
$\left|{[G_{r_{N-1},r_{N};\bar{\psi}}^{n+1}\psi_{0}](s,x)-[G_{r_{N-1},r_{N};\bar{\psi}}^{n+1}\psi_{0}](s,\tilde{x})}\right|\leq\left(c^{\prime}_{N}{\mathrm{e}}^{L_{f}T}+\frac{1+c_{n}(c^{\prime}_{N})}{2}\right)\lVert{x-\tilde{x}}\rVert_{s}$
[…]
for all $s\in[s_{N-1,M-2},s_{N-1,M-1}]$, $x$, $\tilde{x}\in\Omega$. By Theorem 5.4 (ii), letting $n\to\infty$ yields (5.7) on $[s_{N-1,M-2},s_{N-1,M-1}]\times\Omega\times\Omega$ with $12c^{\prime}_{N}{\mathrm{e}}^{L_{f}T}+$ 1 in place of $L_{\bar{\psi}}$.
Step 3. Repeating the procedure of Step 2 finitely many times and making the obvious adjustments every time we pass to a different interval $[r_{k-1},r_{k}]$, we obtain the existence of some constant $L_{\bar{\psi}}\in{\mathbb{R}}_{+}$ such that (5.7) holds on $[0,T]\times\Omega\times\Omega$. ∎
Appendix C On regularity of certain functions on Skorokhod space
Example C.1.
This example is closely related to Example 2.15 (i) in [Keller16SPA]. Let $d=1$. Consider the function $u:[0,T]\times\Omega\to{\mathbb{R}}$ defined by $u(t,x):=\lVert{x(\cdot\wedge t)}\rVert_{\infty}$. Fix $t_{0}\in[0,T)$ and consider the path $x_{0}:=(-2).{\mathbf{1}}_{[t_{0},T)}$.
(i) Note that $u(t_{0},x_{0}\otimes_{t_{0}}(-1))=1$, but for every sufficiently small ${\varepsilon}>0$, $u(t_{0}+{\varepsilon},x_{0}\otimes_{t_{0}+{\varepsilon}}(-1))=2$, i.e., the map $t\mapsto u(t,x_{0}\otimes_{t}(-1))$ is not right-continuous at $t=t_{0}$.
(ii) Note that $-u$ is continuous, but $(t,x)\mapsto-u(t,x\otimes_{t}(-1))$ is not lower semi-continuous as, by (i),
$\liminf_{n\to\infty}\left[-u(t_{0}+n^{-1},x_{0}\otimes_{t_{0}+n^{-1}}(-1))\right]=-2\not\geq-1=-u(t_{0},x_{0}\otimes_{t_{0}}(-1))$
and $\mathbf{d}_{\infty}((t_{0}+n^{-1},x_{0}),(t_{0},x_{0}))=n^{-1}\to 0$ as $n\to\infty$.
Remark C.2.
Continuity (or semi-continuity) of a function $u:[0,T]\times\Omega\to{\mathbb{R}}$ does not, in general, imply continuity (or semi-continuity) of the functions $(t,x)\mapsto u(t,x\otimes_{t}e)$, $e\in{\mathbb{R}}^{d}$ (see Example C.1 as well as (an appropriate modification of) Example 2.15 (iv) in [Keller16SPA]). Also note that this topic (in a very similar but not identical context) has been addressed in section 2.4 of [Keller16SPA].